### v2

# Think How to Think: Mitigating Overthinking with Autonomous Difficulty Cognition in Large Reasoning Models

Abstract
Recent Large Reasoning Models (LRMs) excel at complex reasoning tasks but often suffer from overthinking, generating overly long and redundant reasoning trajectories. To explore its essence, our empirical analysis reveals that LRMs are primarily limited to recognizing task properties (i.e., difficulty levels) like humans before solving the problem, leading to a one-size-fits-all reasoning process. Inspired by this, a pressing and natural question emerges: Can we explicitly bootstrap such ability to alleviate overthinking in LRMs? In this paper, we propose Think-How-to-Think (TH2T), a novel two-stage fine-tuning strategy that progressively inspires LRMs’ difficulty cognition and redundancy cognition of LRMs. Specifically, we first inject difficulty hypnosis into output prefixes to guide the model toward adaptive reasoning depth, trained on a hybrid dataset mixing short and long reasoning paths. Then, we incorporate redundancy hypnosis, which supervises the intermediate reasoning steps to identify and eliminate unnecessary reasoning patterns. Experiments on 7B/14B/32B models demonstrate that TH2T significantly reduces inference costs by over 70% on easy tasks and 40% on hard tasks while maintaining performance stability. The resulting outputs exhibit clear signs of difficulty-aware capabilities and reduced redundancy (e.g., reflection and looping).
Introduction
The emergence of reasoning-focused Large Language Models, also known as Large Reasoning Models (LRMs), has led to substantial improvements in problem-solving capabilities (Sun et al. 2024b; Yu et al. 2024; Team et al. 2025), particularly through the use of elaborate and sequential reasoning enabled by test-time scaling (Snell et al. 2025). In addition to OpenAI-O1 (OpenAI et al. 2024a), other models such as the open-source Deepseek-R1 (Guo et al. 2025) and QwQ (Team 2025) leverage long Chain-of-Thought (Wei et al. 2022; Wang et al. 2023) and adopt reinforcement learning (RL) to approach complex tasks step-by-step, enabling a human-like thinking pattern with reflection, backtracking, and self-validation (Guo et al. 2025; Xie et al. 2025; Liu et al. 2025). This reasoning paradigm demonstrably enhances the models’ problem-solving capabilities and yields promising results. However, lengthy responses introduce the problem of “overthinking” (Team et al. 2025; Chen et al. 2025), wherein model performance initially improves with extended reasoning chains but deteriorates beyond a certain length. For instance, even for trivial queries such as “What is 1 plus 1?”, current models often misjudge the difficulty and produce unnecessarily long outputs—sometimes spanning hundreds of tokens. This tendency toward redundant reasoning not only increases computational cost and latency but also degrades user experience. These observations prompt a critical question: do LRMs possess the capacity to recognize and adapt to varying levels of problem difficulty autonomously?
To answer this question, we start from the perspective of cognitive science, specifically the dual-process theory of human reasoning. This theory posits that human cognition operates via two distinct but complementary systems: System 1, which is fast, concise, and intuitive, and System 2, which is slow, deliberate, and analytical (Kahneman 2011; Evans and Stanovich 2013). Before engaging in problem-solving, humans often perform an initial assessment of a task’s difficulty, allowing them to allocate cognitive resources appropriately between these two systems (Efklides 2011; Flavell 1979). For straightforward tasks, reliance on System 1 suffices, whereas complex problems necessitate the engagement of System 2. This metacognitive ability to evaluate task difficulty and adapt reasoning strategies accordingly, however, is lacking in current reasoning models. As shown in Fig. 2, DeepSeek-R1-Distill-Qwen2.5-7B frequently conflates easy datasets (e.g., GSM8K (Cobbe et al. 2021)) and hard datasets (e.g., AIME2024), treating them with a uniform level of reasoning complexity. These limitations highlight the need for LRMs to develop autonomous difficulty cognition—to “think how to think”.
Current research primarily emphasizes unified CoT compression based on length, yet length is not an accurate proxy for difficulty. In this paper, we propose Think-How-to-Think (TH2T), a novel two-stage fine-tuning strategy that progressively inspires LRMs difficulty cognition and redundancy cognition. In contrast to prompt-based input guidance for instruction adherence, our approach emphasizes output-level interventions, directly intervening inside the reasoning trajectory to establish a self-hypnosis mechanism. Specifically, in the first stage, to encourage the model to develop native difficulty cognition autonomously, we design difficulty-awareness hypnosis and inject it into response prefixes to establish cognitive hints. Combined with a heterogeneous dataset, the trained model enhances its sensitivity to task difficulty, thereby reducing overthinking in simpler tasks through structured cognitive self-guidance.
In the second stage, we conduct fine-grained in-progress redundancy management, enhancing LRM’s redundancy cognition by targeting redundant yet logically complete inference structures (e.g., reflections or double-checks) and terminal reasoning loops commonly observed in complex task generations. The identified structures are injected with redundancy hypnosis while truncating subsequent superfluous content. Overall, our contributions can be outlined as follows:
We observe that current LRMs fall short of perceiving task complexity. To mitigate overthinking, we propose enhancing LRMs with the ability to cognize and adapt to varying levels of problem difficulty autonomously.
To align with human System 1 and System 2 thinking patterns, we propose a novel two-stage fine-tuning strategy that progressively inspires LRMs’ difficulty and redundancy cognition with a constructed heterogeneous dataset with inserted self-hypnosis.
Experiments demonstrate over 70% length reduction on easy tasks and 40% on hard tasks with maintained performance stability, exhibiting clear difficulty-aware capabilities and reduced redundancy.
Literature Review
The advent of LRM (OpenAI et al. 2024a; Guo et al. 2025; Team 2025) enables breakthroughs in problem solving (Sun et al. 2024b; Yu et al. 2024) with Chain-of-Thought (Wei et al. 2022; Wang et al. 2023) and test-time scaling (Snell et al. 2025). Nevertheless, this introduces the challenge of overthinking (Chen et al. 2025; Team et al. 2025). Works of efficient reasoning (Sui et al. 2025) as summarized:
Model-based methods.
Reinforcement Learning motivates some existing efficient reasoning works (Arora and Zanette 2025; Team et al. 2025; Aggarwal and Welleck 2025) to incorporate length penalties. O1-pruner (Luo et al. 2025b) introduces the Length-Harmonizing Reward, integrated with a PPO-style loss. DAST (Shen et al. 2025b) constructs a length-preference dataset and employs SimPO (Meng and Xia 2024). AdaR1 (Luo et al. 2025a) constructs a hybrid reasoning model by merging long and short CoT models and is optimized with DPO (Rafailov et al. 2023). Apart from RL, which is vulnerable and expensive, fine-tuning is another effective post-training strategy (Yang et al. 2025b). CoT-Value (Ma et al. 2025b) regulates reasoning length by a controllable direction in parameter space, enabling dynamic length control. C3oT (Kang et al. 2025) employs GPT-4 as a compressor while retaining key information. TokenSkip (Xia et al. 2025) prunes less important tokens and fine-tune the model to skip the redundant tokens.
Output-based methods.
Implicit-based methods (Deng, Choi, and Shieber 2024; Hao et al. 2024; Saunshi et al. 2025; Cheng and Durme 2024; Shen et al. 2025a) propose to internalize reasoning steps within the latent representation, transitioning reasoning from language space to hidden space. (Saunshi et al. 2025) simulates multiple steps of CoT through a looped transformer layer, increasing the depth of reasoning without additional parameters. However, these implicit techniques introduce performance decline and affect interpretability. There are also dynamic criteria. For example, (Sun et al. 2024a; Xie et al. 2023; Liao et al. 2025) optimize Best-of-N (BoN) decoding with speculation and rejection. (Li et al. 2024; Yang et al. 2025a; Wang et al. 2025) explores possible early exit point. NoThinking (Ma et al. 2025a) prompts reasoning models to directly output final answers without thinking, but heavily relies on sampling, which adversely increases overall overhead.
Prompt-based methods.
Some works (Aytes, Baek, and Hwang 2025; Ding et al. 2024; Ning et al. 2024; Han et al. 2025; Xu et al. 2025; Lee, Che, and Peng 2025) instruct LRMs by various prompts to respond concisely. Specifically, (Renze and Guven 2024) explicitly prompts LLMs to perform concise step-by-step reasoning. SoT (Ning et al. 2024) combines reasoning paradigms with linguistic constraints to minimize token usage. Token-Budget (Han et al. 2025) establishes a token budget estimated by LLM itself, and incorporates this estimation as a token number constraint. (Ge et al. 2025) regulates thinking token distribution by few-shot in-context learning. Nevertheless, prompting relies on precise instruction following and robust task cognition, which will be proven unreliable later.
Overall, most existing works treat efficient reasoning as a unified length compression paradigm. There is a necessity that has been widely overlooked that LRM may need the capacity to autonomously recognize and adapt to varying levels of problem difficulty, which aligns with System 1 and System 2 thinking patterns of human cognitive theory.
Preliminary and Observation
Background and Notations
LLMs.
Given a prompted question , conventional LLM generates a response by iteratively predicting each response token conditioned on and generated tokens . This is termed as System 1 (Li et al. 2025b) thinking:
where [Prompt] and [Answer] are question and response , which is often short and concise.
LRMs.
Whereas in reasoning-focused LRMs, the generation process is split into two stages, e.g., thinking and conclusion, which is referred to as System 2 (Li et al. 2025b):
where and are two delimiters denoting the start and end of reasoning, [Thoughts] and [Conclusion] are detailed reasoning process and concise answer, respectively. Formally, the reasoning part [Thoughts] reveals a trajectory of problem solving, known as CoT (Wei et al. 2022; Wang et al. 2023). At the micro level, some temporally consecutive tokens in [Thoughts] can be aggregated into short chunk based on semantics, logic and grammar:
Similarly, each reasoning chunk in this paradigm is autoregressively generated by conditioning on question and previous reasoning thoughts: . This explicit decomposition not only improves interpretability through transparent exposure of reasoning steps but also enhances the model’s capability for complex tasks. Specifically, some chunks are well structured by prefix keywords such as “Wait”, “However”, “Alternatively” (Guo et al. 2025), marking the subsequent chunks will undergo logical backtracking, self-verification, or exploration. Works (Li et al. 2025a) have shown that reflection structures are crucial to LRMs. Under this setting, the “overthinking” phenomenon is that many seem delicate yet contribute nothing to reasoning logic, resulting in prominent overhead.
Observation: Reasoning Process Intervention
To address the overthinking issue, a straightforward approach is difficulty reminders—encouraging LRMs to self-regulate output length based on task difficulty judgment through prompts like “Please think quickly if you find it easy”. However, this conventional interventional technique necessitates two prerequisites: accurate instruction following and robust task difficulty cognition. While prior studies (Wu et al. 2025; Han et al. 2025) have highlighted the inherent challenges of the former, we have already shown that current LRMs struggle to recognize varying levels of problem difficulty in Fig. 2.
Instead of using a difficulty reminder in the prompt, we claim that a promising solution is transferring the intervention from the prompt to the inner reasoning process. We conduct a pre-experiment in Fig. 3, pretending the model has already correctly recognized the difficulty level of samples. Specifically, for a simple problem, we change the model input from “User: [Prompt]+[Question]. Assistant:” to “User: [Prompt]+[Question]. Assistant: This is a simple question, let’s think quickly.”. While for hard problems, the corresponding predefined response prefix is “It seems difficult, let’s think thoroughly.”. Fig. 3 shows that predefined difficulty awareness indeed helps to reduce response length, especially matched awareness (“simple, quickly” for easy and “difficult, thoroughly” for hard is matched, the reverse is unmatched).
We define them as self-hypnosis that acts as a trigger in the output, guiding the subsequent inference trajectory. In fact, the reflection structures in LRMs (Guo et al. 2025; Li et al. 2025a) are essentially similar triggers aiming to intervene in subsequent think chunks. Compared to explicit prompts, we argue that these built-in triggers offer three key advantages: (1) native flexibility, (2) robustness, and (3) superior instruction-following capabilities. In the next section, we introduce how we guide LRMs to generate native interventional triggers of difficulty and redundancy cognition, autonomously.
Methodology
In this section, we describe our two-stage fine-tuning strategy Think-How-to-Think (TH2T) for mitigating overthinking, as shown in Fig. 4.
Stage 1: Difficulty Cognition Injection
We aim to endow LRM with the capability to assess the difficulty of a given problem. We define two datasets: for easy problems, and for difficult ones. Two models from the same series were selected to ensure consistency in response: for basic LLM with prompted Short-CoT capability (e.g., Qwen2.5 series) and for advanced LRM (e.g., DeepSeek-R1-Distill-Qwen2.5 series). For each problem in , the responses are recorded as:
where means the response of on is correct. can replicate the majority correct responses of on , which means: , wherein significant additional overhead yields limited performance gain. This motivates us to preserve ’s potential for efficiency on with satisfactory performance. Our desired thinking pattern of is responsed concisely on and elaborately on . To align with this objective, we construct a draft dataset grounded in difficulty differentiation:
where , similar to Eq. (5). aims to maintain the performance gain of on .
To enable to perceive problem difficulty level autonomously, we introduce the trials of Fig. 3. We inject specific difficulty-aware cognitive Hypnosis into the prefix of answer within , which, as demonstrated before, functions as interventional triggers guiding the reasoning trajectory:
If , the difficulty hypnosis H is defined as follows: “hypnosis This is a simple question, let’s think quickly./hypnosis”. If , then “hypnosisIt seems difficult, let’s think thoroughly./hypnosis”. The markers “hypnosis” and “/hypnosis” are defined as indicators for inserted difficulty hypnosis. The injected hypnosis is recognized as an embedded difficulty label to train to generate native difficulty cognition on the refined dataset , stimulating a concise thinking pattern on easy tasks. The execution of stage 1 is outlined at the upper section of the algorithmic framework Fig. (4).
Stage 2: Redundancy Cognition Injection
We extend difficulty-hypnosis of Stage 1 to redundancy-hypnosis, guiding to identify redundant structures within the reasoning process on hard tasks.
[…]
where is the difficulty-hypnosis introduced in Stage 1, are valid reasoning chunks to be kept, is the unnecessary reflection structure inside the reasoning process without effective contribution to the reasoning process. These chunks often carry out backtracking, self-verification, or exploration, but unfortunately, are nothing more than superfluous overhead. is a similar reflection structure; nonetheless, it primarily manifests in the final phase of the reasoning process, continuously iterating until attains the maximum output length.
We define a Determinator to judge whether some is logically equivalent to or (i.e., redundant or looping). If yes, we inject redundancy hypnosis and truncate subsequent content. In case of , original is refined as:
wherein the redundancy hypnosis is defined as follows: “hypnosis Everything seems ok, let’s move on./hypnosis”. We use GPT-4 (OpenAI et al. 2024b) to implement the determinator.
In case of , on one hand, the training set lacks a sufficient number of loop structures; on the other hand, excessively deferred looping hinders the generation of shorter-length samples. Thus, we select intermediate chunk with words like “Wait”, “However”, “Alternatively” and iterate it as a simulation of looping, followed by hypnosis injection and truncation:
wherein the looping hypnosis is defined as follows: “hypnosis Oh, I’m stuck in a loop. Time to break out. /hypnosis”. The choice of iterate number can be random, and we also try a probe-like approach (Zhang et al. 2025). We use DeepSeek-R1 to identify chunks before which the prior content can already induce the correct answer, referred to as completeness requirements. The last-layer hidden states of the selected chunk tail is extracted to train a simple two-layer MLP as Determinator. Experiments indicate it yields similar results to the random approach, so we prefer the latter.
Notably, we adopt truncation instead of mere substitution, as truncation allows for the construction of shorter samples. Meanwhile, generation continuity is ensured by remaining untruncated samples in the dataset, since we only apply Stage 2 to part of the samples in . Stage 2 execution is outlined in Fig. 4.
Experiments
Experiment Setup
Models.
Dataset Construction.
For dataset construction, we mainly adopt mathematical reasoning tasks, which are friendly for difficulty quantization. In stage 1, the GSM8K and MATH datasets are combined after filtering out errors and injecting difficulty hypnosis. In stage 2, redundancy and looping hypnosis are injected into difficult samples. The composition ratio of this heterogeneous dataset is 2:1:1.
Baselines.
We compare TH2T with two prompt-based methods, i.e., D-Prompt and NoThinking (Ma et al. 2025a), and two fine-tuning methods, i.e., Tokenskip (Xia et al. 2025), CoT-Valve (Ma et al. 2025b). D-Prompt directly promotes LRM to regulate its output length based on self-judgment of task difficulty, through difficulty reminder in the prompt, as discussed in Preliminary Part. The remaining methods are from recent publications.
| Methods | GSM8K | MATH-500 | ||||||||||
| Acc | Gain | Len | Reduc | Latency | SpdUp | Acc | Gain | Len | Reduc | Latency | SpdUp | |
| DeepSeek-R1-Distill-Qwen-7B | ||||||||||||
| Original | 91.51 | 1057 | 20.29 | 85.2 | 2857 | 87.69 | ||||||
| [1pt/2pt] D-Prompt | 90.59 | 0.92 | 906 | 14.3% | 16.38 | 1.24 | 86.0 | 0.8 | 2726 | 4.6% | 81.54 | 1.08 |
| NoThinking | 88.47 | 3.04 | 243 | 76.9% | 3.34 | 6.07 | 79.4 | 5.8 | 699 | 75.5% | 11.81 | 6.43 |
| [1pt/2pt] TokenSkip | 88.01 | 3.50 | 638 | 39.6% | 10.64 | 1.91 | 80.2 | 5.0 | 2109 | 26.2% | 53.94 | 1.62 |
| CoT-Valve | 87.90 | 3.61 | 235 | 77.8% | 3.31 | 6.13 | 78.6 | 6.6 | 747 | 73.9% | 12.31 | 7.12 |
| (w/o D.H.) | 90.83 | 0.68 | 503 | 47.4% | 7.80 | 2.60 | 85.2 | 0.0 | 1906 | 33.2% | 46.68 | 1.88 |
| (w/o R.H.) | 91.96 | 0.48 | 397 | 62.4% | 5.99 | 3.39 | 86.0 | 0.8 | 2265 | 20.7% | 60.22 | 1.46 |
| TH2T | 92.87 | 1.36 | 275 | 74.0% | 3.88 | 5.23 | 86.8 | 1.6 | 1772 | 38.0% | 41.65 | 2.11 |
| DeepSeek-R1-Distill-Qwen-14B | ||||||||||||
| Original | 93.63 | 584 | 9.38 | 86.6 | 2306 | 62.65 | ||||||
| [1pt/2pt] D-Prompt | 93.91 | 0.28 | 606 | 3.9% | 9.95 | 0.94 | 86.6 | 0.0 | 2297 | 0.3% | 62.60 | 1.01 |
| NoThinking | 88.41 | 3.48 | 206 | 64.9% | 3.06 | 3.07 | 76.0 | 10.6 | 612 | 73.5% | 10.05 | 6.23 |
| [1pt/2pt] TokenSkip | 89.30 | 4.33 | 314 | 46.2% | 4.79 | 1.96 | 73.2 | 13.4 | 1356 | 41.2% | 28.11 | 2.23 |
| CoT-Valve | 92.70 | 0.90 | 526 | 9.9% | 8.13 | 1.15 | 84.2 | 2.4 | 2111 | 8.4% | 53.95 | 1.16 |
| (w/o D.H.) | 93.69 | 0.06 | 338 | 42.2% | 4.82 | 1.95 | 86.2 | 0.4 | 1549 | 32.8% | 33.93 | 1.85 |
| (w/o R.H.) | 93.47 | 0.16 | 299 | 51.2% | 4.34 | 2.16 | 85.8 | 0.8 | 1706 | 26.0% | 39.83 | 1.57 |
| TH2T | 94.08 | 0.45 | 246 | 58.1% | 3.38 | 2.78 | 86.4 | 0.2 | 1343 | 41.8% | 27.84 | 2.25 |
| DeepSeek-R1-Distill-Qwen-32B | ||||||||||||
| Original | 95.60 | 717 | 12.14 | 87.2 | 2357 | 63.02 | ||||||
| [1pt/2pt] D-Prompt | 95.60 | 0.00 | 663 | 7.6% | 11.12 | 1.09 | 86.8 | 0.4 | 2198 | 6.8% | 57.11 | 1.10 |
| NoThinking | 93.86 | 1.74 | 228 | 68.2% | 3.25 | 3.74 | 80.2 | 7.0 | 659 | 72.0% | 10.98 | 5.74 |
| [1pt/2pt] TokenSkip | 96.54 | 0.94 | 423 | 41.0% | 6.28 | 1.93 | 79.8 | 7.4 | 1567 | 33.5% | 34.26 | 1.84 |
| CoT-Valve | 96.10 | 0.50 | 781 | 8.9% | 12.56 | 0.97 | 85.3 | 1.9 | 2263 | 3.9% | 60.21 | 1.05 |
| (w/o D.H.) | 96.08 | 0.48 | 361 | 49.7% | 5.47 | 2.22 | 87.4 | 0.2 | 1863 | 20.9% | 45.18 | 1.39 |
| (w/o R.H.) | 95.75 | 0.15 | 323 | 55.0% | 4.84 | 2.51 | 86.8 | 0.4 | 2114 | 10.3% | 53.99 | 1.17 |
| TH2T | 97.04 | 1.44 | 263 | 63.3% | 3.51 | 3.46 | 88.0 | 0.8 | 1733 | 26.5% | 40.13 | 1.57 |
| Bench | Difficulty Level | Cognition | |
| Original | TH2T | ||
| GSM8K | all easy | 14.9% | 98.3% |
| MATH | lv.1-2 (easy) | 9.8% | 93.4% |
| lv.4-5 (hard) | 12.0% | 91.5% | |
| AIME2024 | all hard | 16.7% | 96.7% |
| OmniMath | lv.1-2 (easy) | 20.0% | 96.7% |
| lv.4-5 (hard) | 13.3% | 93.3% | |
Evaluations.
We evaluate on 4 mathematical benchmarks (GSM8K, MATH, AIME2024 and Omni-MATH) and 1 QA benchmark GPQA (Rein et al. 2024). Metrics include accuracy, average number of output tokens, and inference latency per sample. We also report the success rate of the difficulty awareness, redundancy, and looping optimization.
The detailed experiment settings are in the appendix.
Experimental Results
Main Results.
The performance of TH2T on GSM8K and MATH is illustrated in Tab. 1. Key observations include: 1) TH2T exhibits minimal accuracy degradation; more precisely, most settings witness slight improvement. 2) TH2T achieves the best reduction in reasoning length among methods with stable accuracy—up to 74% on GSM8K and 38% on MATH with the 7B model, translating to over 5 and 2 latency speedups, respectively. 3) Greater gains on easier tasks, validating TH2T’s ability to adapt reasoning effort based on task difficulty. 4) These observations are consistent across 7B, 14B and 32B models. As shown in Fig. 1, other methods (e.g., NoThinking and CoT-Value) achieve comparable compression ratios in token length but at the cost of severe accuracy degradation (often 10 points), underscoring TH2T’s robustness in balancing efficiency and reliability.
Difficulty Cognition.
Does TH2T present better difficulty cognition?. As shown in Tab. 2, the entirety of GSM8K, MATH lv.1-2, and Omni-Math lv.1-2 are set as “Easy” while the remaining are set as “Hard”. The original R1-Distilled-7B model is prompted to predict its own understanding of problem difficulty and choose between “Easy”, “Medium”, and “Hard”. TH2T can response with native self-hypnosis in the prefix like “This is a simple question” or “It seems difficult”, which indicates the cognition of task difficulty as “Easy” or “Hard”. As shown in Tab. 2 and Fig. 2, the original LRM demonstrates ambiguous recognition of task difficulty, uniformly categorizing tasks of varying complexity as medium level. Whereas, our approach exhibits a clear distinction with high precision over 90+%, validating the effectiveness of the first stage of TH2T.
Redundancy Cognition.
Does TH2T present better redundancy cognition? Tab. 3 shows the average number of reflective and looping structures. Comparison is made between R1-Distilled-7B and Ours-7B. We observe that: 1) TH2T effectively suppresses the number of reflection chunks by 10 on easy tasks and 3 on hard tasks, suggesting that extensive reflection is unnecessary for strong reasoning. 2) The looping problem in TH2T undergoes substantial restraint. Notably, 59.5% of incorrect answers by the original model on MATH exhibit tail recursion, which is mitigated to 21.4% in ours. As shown in Fig. 6(b), we observe zero loop structure in GSM8K. This distinct pattern of short-length concentration underscores the efficacy of Stage-2 of TH2T, presenting a more concise reasoning structure.
| Bench | Split | Reflective | Looping | ||
| Original | TH2T | Original | TH2T | ||
| GSM8K | Overall | 10.3 | 1.2 | 2.0% | 0.0% |
| –correct | 4.8 | 0.0 | 0.4% | 0.0% | |
| –incorrect | 69.4 | 10.8 | 22.3% | 0.0% | |
| MATH | Overall | 27.0 | 12.3 | 10.4% | 4.0% |
| –correct | 14.9 | 6.8 | 1.9% | 1.0% | |
| –incorrect | 96.5 | 39.1 | 59.5% | 21.4% | |
Generalization Ability.
To validate the generalization on OOD tasks, we evaluate on 2 more mathematical benchmarks and 1 QA benchmark. As shown in Fig. 5, accuracy maintenance and token length reduction are consistent.
Results of Llama series model are in the appendix.
Ablation Study
Ablation on Stages.
We explore two more settings: 1) without Difficulty Hypnosis in Stage-1, and 2) without Redundancy Hypnosis in Stage-2, which are annotated as w/o D.H. and w/o R.H. in Tab. 1. From the perspective of response reduction, we observe that D.H. introduces more gain on easy task GSM8K than on hard task MATH (e.g., 26.6% vs 4.8%, on 7B model), while R.H. exerts a greater impact on MATH (11.6% vs 17.3%). As we illustrate during method formulation, we design Difficulty Cognition Injection Stage to build native perception on easy tasks and Redundancy Cognition Injection Stage to mitigate non-essential structures that frequently occur in hard tasks. These results demonstrate the rationality of our two-stage strategy.
Ablation on Starting Model.
Instead of training on LRMs, we also try directly starting from a Short-CoT model with the same data in Tab. 4, which reveals a notable failure in performance and response length. This motivates us to include an underlying Stage-0 in Fig. 4 to start from an LRM, instead of teaching a Short-CoT model from scratch.
| Method | GSM8K | MATH-500 | ||
| Accuracy | Length | Accuracy | Length | |
| Qwen2.5-7B | 86.20 | 280 | 62.80 | 629 |
| [1pt/2pt] TH2T | 72.48 | 238 | 66.20 | 2333 |
| R1-Distill-7B | 91.51 | 1057 | 85.2 | 2857 |
| [1pt/2pt] TH2T | 92.87 | 275 | 86.8 | 1772 |
Ablation on Training Steps.
We examine the dynamics of accuracy and average token length during the training process (figure in appendix), and obtain two interesting findings: 1) As training steps increase, accuracy initially declines, then returns to the original level, with occasional positive gains during the process. 2) For hard tasks like MATH, the response length exhibits a reverse increase after exceeding a certain number of steps. This inspires us to set the training steps to 6400 to prevent model degradation.
Conclusion
This paper presents TH2T, a novel two-stage fine-tuning strategy for large reasoning models (LRMs) that effectively addresses the challenge of overthinking. Our empirical analysis reveals that current LRMs are primarily limited to recognizing task difficulty, exhibiting a necessity for aligning with human System 1 and System 2 thinking patterns. Motivated by this, TH2T progressively enables native difficulty and redundancy cognition based on a heterogeneous dataset and proposed self-hypnosis mechanism. To the best of our knowledge, we are the first to inject difficulty/redundancy/looping hypnosises as built-in triggers for reasoning trajectory intervention. Experiments demonstrate that TH2T significantly reduces inference length with maintained performance stability, exhibiting clear difficulty and redundancy sensitivity. Future research can explore finer-grained reasoning trajectory intervention.
Supplementary Materials
More Experimental Details
Dataset Construction
The datasets used in this paper incorporate difficulty level labeling as a key consideration, so we mainly use mathematical reasoning tasks. Our TH2T is a two-stage strategy, with tailored dataset construction for each phase, followed by sequential fine-tuning. 1) During the first stage, we use the GSM8K and MATH training datasets to obtain and , respectively. The former is generally considered simple, whereas the latter is regarded as difficult and is explicitly categorized into difficulty levels ranging from 1 to 5, and we pick samples between lv.3 and lv.5. After filtering out errors, and are balanced at a 1:1 ratio. After injecting difficulty hypnosis , we construct of 6.4K samples. 2) In Stage-2, are injected with redundancy hypnosis and looping hypnosis , forming a final training dataset. We observe that the meticulous composition ratio of the constructed dataset does not significantly affect the final performance. Hence, we combine and in a ratio of 1:1. Meanwhile, the redundancy hypnosis and looping hypnosis are also injected into in a ratio of 1:1. The overall composition is set as 2:1:1.
Baselines
In our main experiments, we compare TH2T with two prompt-based methods and two fine-tuning methods.
D-Prompt directly promotes LRM to regulate its output length based on self-judgment of task difficulty based on Difficulty Reminder in Prompt, as discussed in our Preliminary Part.
NoThinking bypasses the thinking process via prompting for all difficulty-level questions.
Tokenskip constructs a dataset where less important tokens are pruned and fine-tuned to enable the model to skip the redundant tokens in inference, regardless of task difficulty.
CoT-Valve enables dynamic control of CoT length by identifying and leveraging a controllable direction in parameter space, reducing inference cost by compressing reasoning paths based on task difficulty.
We replicate the first two baselines and report the relative performance improvement of the latter two from the raw papers due to variability of the model and implementation methods.
Benchmarks
We evaluate on 4 mathematical benchmarks and 1 QA benchmark. The main results are on the test set of in-distribution GSM8K and MATH. Meanwhile, AIME2024 and Omni-MATH are employed as out-of-distribution test sets. For Omni-MATH, we select 30 samples from difficulty ranges 1-2 and 4-5 as easy/hard instances, while AIME2024 is considered hard. To explore generalization ability in other domains, we include the GPQA benchmark to test performance on QA tasks.
Metrics
Main results are assessed on three metrics: accuracy, average number of output tokens, and inference latency per sample. To substantiate the research objectives, we incorporate the success rate of the difficulty awareness. We also report the redundancy cognition optimization, which includes the reduction of redundant structures (e.g., “Wait”, “Alternatively”, or “Check”), as well as the reflection looping ratio.
Implementation Details
We adopt LoRA, a widespread LLM fine-tuning approach, to train our models with a rank , scaling parameter , and learning rate . All experiments are conducted on 2NVIDIA H800 GPU (80GB). For the decoding strategy, we employ greedy decoding, and the maximum generation length equals 8192 (except 16,384 for AIME2024). The random seed is fixed to 42.
| Methods | GSM8K | MATH-500 | AIME2024 | OmniMath | GPQA | Average | ||||||
| Acc | Len | Acc | Len | Acc | Len | Acc | Len | Acc | Len | Acc | Len | |
| DeepSeek-R1-Distill-Llama-8B | ||||||||||||
| Original | 75.36 | 260 | 61.8 | 701 | 3.3 | 1378 | 26.67 | 844 | 34.34 | 704 | 40.29 | 777 |
| D-Prompt | 75.14 | 261 | 62.2 | 596 | 6.6 | 1100 | 28.33 | 808 | 34.34 | 669 | 41.32 | 687 |
| NoThinking | 75.24 | 257 | 62.2 | 580 | 6.6 | 1019 | 28.33 | 714 | 32.32 | 619 | 40.94 | 638 |
| TH2T | 75.58 | 245 | 61.6 | 498 | 13.3 | 861 | 40.00 | 561 | 40.91 | 526 | 46.28 | 538 |
More Experimental Results
Results on Llama Model
We provide detailed experimental results of the R1-Distilled-Llama-8B model evaluated on 4 mathematical benchmarks (i.e., GSM8K, MATH, AIME2024, and OmniMath) and 1 QA benchmark (i.e., GPQA). We report accuracy and average token length for comparison in Tab. 5.
Acc and Length on Detailed Difficulty Levels
We report a breakdown of the accuracy and average token length distribution across the complete spectrum of difficulty levels on 4 tasks. The comparison is made between original R1-Ditstilled-Qwen–7B and TH2T-7B. Each question in the MATH and OmniMath datasets is labeled with a difficulty level, ranging from 1 to 5 and from 1 to 8, respectively. The higher the numerical level, the more difficult the question. Meanwhile, all questions in GSM8K and AIME2024 are considered easy and very difficult, respectively. Results are illustrated in Fig.7, and we obtain the following observations: 1) Both models generate progressively longer and more elaborate responses as the difficulty level increases. 2) across all difficulty levels in all tasks, TH2T exhibits significant length reduction, as well as stable performance maintenance. 3) The lower the difficulty level, the more significant the reduction in length.
Dynamics of Accuracy and Token Length
We report the dynamics of accuracy and average token length during the training steps in Fig. 8, and obtain two interesting findings: 1) As training steps increase, accuracy initially declines, then returns to the original level, with occasional positive gains during the process. 2) For hard tasks like MATH, the response length exhibits a reverse increase after exceeding a certain number of steps. This inspires us to set the training steps to 6400 to prevent model degradation.
Visualization of the Effect on Token Length
We collect the first 50 samples from GSM8K and MATH test set in sequential order, and compare the output token length distribution between our TH2T and the original DeepSeek-R1-Distill-Qwen-7B in Fig. 9. We can see TH2T significantly reduces response length.
More Explanation on Motivation
We investigate the efficiency issue from a suspicion: Are current LRMs treating various task types equally with identical response strategy? We compare the accuracy gain and inference length scaling ratio between Short-CoT model (Qwen2.5-7B) and Long-CoT model (DeepSeek-R1-Distill-Qwen2.5-7B) on easy GSM8K and hard MATH-500 benchmarks. From Fig. 10, we obtain three findings:
Across the easy and difficult tasks, the current reasoning model shows similar token length scaling ratio (i.e., both around 400%), which verifies our initial suspicion that LRMs are allocating similar additional computing resources to tasks with varying difficulty levels.
Despite incurring minimal costs, basic LLM can still replicate the majority of correct responses (89.9%) of LRM’s on the easy tasks.
Compared with difficult tasks, the easy task only witnesses a minor accuracy gain (5.3% vs 22.4%) under the same computation resource allocation strategy, which is not what we expected. We aim to maintain this marginal performance improvement while keeping the associated token length increase at a minimal level.
These raise a reasonable speculation: LRMs fall short of perceiving task complexity, since massive inference resources on easy ones are superfluous and lead to limited performance gain. As discussed in Results on Redundancy Cognition of the main paper, we observe severe redundant structures in LRM’s responses to hard tasks, including unnecessary reflection along with a remarkably high proportion of reasoning loops in incorrect answers. These findings lead to the two objectives of this paper:
Enabling LRMs to think how to think, i.e., autonomously recognizing and adapting to variant levels of problem difficulty (results and analysis are presented in the Results of Difficulty Cognition section of the main paper).
Mitigating LRM overthinking by promoting redundancy cognition of internal reasoning structure (results and analysis are presented in the Results of Redundancy Cognition section of the main paper).
Prompts
For reproducibility, we provide the complete list of prompts adopted in our experiments.
Difficulty level evaluation
Fig. 11 shows the prompt template we used during difficulty level evaluation.
Baseline
Fig. 12 is the difficulty reminder for the baseline D-Prompt, which is also adopted in the pre-experiment. Fig. 13 is the setting of baseline NoThinking.
Design of Determinator
During Stage 2 of TH2T, we design a Determinator. To identify redundancy point, i.e., , the contribution effectiveness of reflective structures is assessed as shown in Fig. 15 and Fig. 16. Furthermore, when injecting hypnosis, the completeness requirement in Fig. 14 is adopted.
Limitations and Future Work
Our current approach primarily centers on standard fine-tuning using our generated heterogeneous dataset. While this method effectively enhances model performance within our present framework, it may not fully leverage the potential for optimized efficiency. Advanced reinforcement learning (RL)-based training strategies may offer a promising avenue for further performance improvements. However, RL can be very unstable and computationally intensive. Due to computational constraints, we defer the exploration of sophisticated RL techniques to future research.
We also recognize that theoretically, models of different series and parameter scales have distinct capability boundaries, leading to varying levels of task difficulty perception. In this phase, we introduce a common difficulty evaluation standard applicable to models at different scales. We plan to explore difficulty definition methods tailored to models of variant capability ranges in the future, aiming to advance more autonomous difficulty adaptation and perception strategies. One possible solution is to depend on the correctness ratio during sampling. However, a significant limitation of this method lies in its substantially increased computational overhead.
Case Study

### v3

@@
 Recent Large Reasoning Models (LRMs) excel at complex reasoning tasks but often suffer from overthinking, generating overly long and redundant reasoning trajectories.
-To explore its essence, our empirical analysis reveals that LRMs are primarily limited to recognizing task properties (i.e., difficulty levels) like humans before solving the problem, leading to a one-size-fits-all reasoning process.
-Inspired by this, a pressing and natural question emerges: Can we explicitly bootstrap such ability to alleviate overthinking in LRMs?
-In this paper, we propose Think-How-to-Think (TH2T), a novel two-stage fine-tuning strategy that progressively inspires LRMs’ difficulty cognition and redundancy cognition of LRMs.
-Specifically, we first inject difficulty hypnosis into output prefixes to guide the model toward adaptive reasoning depth, trained on a hybrid dataset mixing short and long reasoning paths.
-Then, we incorporate redundancy hypnosis, which supervises the intermediate reasoning steps to identify and eliminate unnecessary reasoning patterns.
-Experiments on 7B/14B/32B models demonstrate that TH2T significantly reduces inference costs by over 70% on easy tasks and 40% on hard tasks while maintaining performance stability.
-The resulting outputs exhibit clear signs of difficulty-aware capabilities and reduced redundancy (e.g., reflection and looping).
-Introduction
-The emergence of reasoning-focused Large Language Models, also known as Large Reasoning Models (LRMs), has led to substantial improvements in problem-solving capabilities (Sun et al.
-2024b; Yu et al.
-2024; Team et al.
-2025), particularly through the use of elaborate and sequential reasoning enabled by test-time scaling (Snell et al.
-2025).
-In addition to OpenAI-O1 (OpenAI et al.
-2024a), other models such as the open-source Deepseek-R1 (Guo et al.
-2025) and QwQ (Team 2025) leverage long Chain-of-Thought (Wei et al.
-2022; Wang et al.
-2023) and adopt reinforcement learning (RL) to approach complex tasks step-by-step, enabling a human-like thinking pattern with reflection, backtracking, and self-validation (Guo et al.
-2025; Xie et al.
-2025; Liu et al.
-2025).
+To explore its essence, our empirical analysis reveals that LRMs are primarily limited to recognizing task properties (i.e., difficulty levels) like humans before solving the problem, leading to a one-size-fits-all reasoning strategy.
+This observation motivates a fundamental question: Can we explicitly bootstrap such ability to alleviate overthinking in LRMs?
+To this end, we propose Think-How-to-Think (TH2T), a novel two-stage fine-tuning strategy that progressively inspires LRMs’ difficulty cognition and redundancy cognition of LRMs.
+Specifically, we first inject Difficulty Dypnosis into output prefixes as cues for global, prospective reasoning strategy selection, stimulating the model’s sharper sensitivity to task complexity and adaptive control of reasoning depth.
+Then, we incorporate Redundancy Hypnosis into in-progress reasoning steps, which serve as local, retrospective signals for behavior correction by identifying and eliminating superfluous reasoning detours.
+Experiments across 7B/14B/32B models demonstrate that TH2T significantly reduces inference costs by over 70% on easy tasks and 40% on complex ones without compromising performance.
+The resultant models exhibit a nascent ability for difficulty-aware reasoning, effectively mitigating behaviors like excessive reflection and looping, thereby paving the way for more cognitively efficient LRMs.
+Think How to Think: Mitigating Overthinking with Autonomous Difficulty Cognition in Large Reasoning Models
+Yongjiang Liu1, Haoxi Li1, Xiaosong Ma2, Jie Zhang1, Song Guo1, 1The Hong Kong University of Science and Technology, 2The Hong Kong Polytechnic University
+1 Introduction
+The emergence of Large Reasoning Models (LRMs) has led to substantial improvements in problem-solving capabilities Sun et al.
+(2024b); Yu et al.
+(2024); Team et al.
+(2025), particularly through the use of elaborate and sequential reasoning enabled by test-time scaling Snell et al.
+(2025).
+Many models, such as OpenAI-O1 OpenAI et al.
+(2024a), Deepseek-R1 Guo et al.
+(2025), QwQ Team (2025) leverage long Chain-of-Thought Wei et al.
+(2022); Wang et al.
+(2023) to approach complex tasks step-by-step, enabling a human-like thinking pattern with reflection, backtracking, and self-validation Xie et al.
+(2025); Liu et al.
+(2025).
 This reasoning paradigm demonstrably enhances the models’ problem-solving capabilities and yields promising results.
-However, lengthy responses introduce the problem of “overthinking” (Team et al.
-2025; Chen et al.
-2025), wherein model performance initially improves with extended reasoning chains but deteriorates beyond a certain length.
-For instance, even for trivial queries such as “What is 1 plus 1?”, current models often misjudge the difficulty and produce unnecessarily long outputs—sometimes spanning hundreds of tokens.
-This tendency toward redundant reasoning not only increases computational cost and latency but also degrades user experience.
-These observations prompt a critical question: do LRMs possess the capacity to recognize and adapt to varying levels of problem difficulty autonomously?
-To answer this question, we start from the perspective of cognitive science, specifically the dual-process theory of human reasoning.
-This theory posits that human cognition operates via two distinct but complementary systems: System 1, which is fast, concise, and intuitive, and System 2, which is slow, deliberate, and analytical (Kahneman 2011; Evans and Stanovich 2013).
-Before engaging in problem-solving, humans often perform an initial assessment of a task’s difficulty, allowing them to allocate cognitive resources appropriately between these two systems (Efklides 2011; Flavell 1979).
+However, lengthy responses introduce the problem of “overthinking” Chen et al.
+(2025), wherein model performance initially improves with extended reasoning chains but deteriorates beyond a certain length.
+This inefficiency is starkly evident when trivial queries, such as “What is 1 plus 1?”, trigger disproportionately complex reasoning paths.
+Such redundant deliberation not only inflates computational cost and latency but also degrades the user experience.
+These lead us to a central inquiry: Do LRMs possess the cognitive judgment to dynamically regulate their reasoning depth in response to varying problem difficulty?
+Inspiration for a solution lies in the human cognitive science: the dual-process theory of human reasoning.
+This theory posits a cognitive architecture composed of two distinct but complementary systems: a fast, concise, and intuitive System 1, and a slow, deliberate, and analytical System 2 Kahneman (2011); Evans and Stanovich (2013).
+Before engaging in problem-solving, humans often perform an initial assessment of a task’s difficulty, allowing them to allocate cognitive resources appropriately between these two systems Efklides (2011); Flavell (1979).
 For straightforward tasks, reliance on System 1 suffices, whereas complex problems necessitate the engagement of System 2.
-This metacognitive ability to evaluate task difficulty and adapt reasoning strategies accordingly, however, is lacking in current reasoning models.
-As shown in Fig.
-2, DeepSeek-R1-Distill-Qwen2.5-7B frequently conflates easy datasets (e.g., GSM8K (Cobbe et al.
-2021)) and hard datasets (e.g., AIME2024), treating them with a uniform level of reasoning complexity.
+Our empirical analysis (Fig.
+2) reveals that they approach disparate tasks—from the straightforward (e.g., GSM8K) to the highly complex (e.g., AIME2024)—with a monolithic and inefficient reasoning strategy.
 These limitations highlight the need for LRMs to develop autonomous difficulty cognition—to “think how to think”.
-Current research primarily emphasizes unified CoT compression based on length, yet length is not an accurate proxy for difficulty.
 In this paper, we propose Think-How-to-Think (TH2T), a novel two-stage fine-tuning strategy that progressively inspires LRMs difficulty cognition and redundancy cognition.
 In contrast to prompt-based input guidance for instruction adherence, our approach emphasizes output-level interventions, directly intervening inside the reasoning trajectory to establish a self-hypnosis mechanism.
-Specifically, in the first stage, to encourage the model to develop native difficulty cognition autonomously, we design difficulty-awareness hypnosis and inject it into response prefixes to establish cognitive hints.
-Combined with a heterogeneous dataset, the trained model enhances its sensitivity to task difficulty, thereby reducing overthinking in simpler tasks through structured cognitive self-guidance.
-In the second stage, we conduct fine-grained in-progress redundancy management, enhancing LRM’s redundancy cognition by targeting redundant yet logically complete inference structures (e.g., reflections or double-checks) and terminal reasoning loops commonly observed in complex task generations.
-The identified structures are injected with redundancy hypnosis while truncating subsequent superfluous content.
+In the first stage, we design Difficulty Hypnosis (DH) to stimulate the model’s awareness of task difficulty.
+DH introduces difficulty cues into output prefixes, which act as guiding signals for global, prospective strategy selection.
+With DH supervision from a hybrid dataset featuring both concise and elaborate reasoning trajectories, the model calibrates its sensitivity to task difficulty and adaptive control of reasoning depth, ultimately curtailing the impulse to overthink on straightforward problems through structured cognitive self-guidance.In the second stage, we design Redundancy Hypnosis (RH) to cultivate the model’s awareness of redundant reasoning.
+RH specifically targets fine-grained inference patterns such as redundant yet logically complete structures (e.g., reflections or double-checks) and terminal reasoning loops frequently observed in complex task generation.
+By embedding redundancy cues into these in-progress structures, RH provides local, retrospective signals for behavior correction, enabling the model to skip or truncate superfluous steps.
+In this way, the model strengthens its redundancy cognition and reduces superfluous reasoning detours.
 Overall, our contributions can be outlined as follows:
-We observe that current LRMs fall short of perceiving task complexity.
+We reveal that current LRMs fall short of perceiving task complexity.
 To mitigate overthinking, we propose enhancing LRMs with the ability to cognize and adapt to varying levels of problem difficulty autonomously.
-To align with human System 1 and System 2 thinking patterns, we propose a novel two-stage fine-tuning strategy that progressively inspires LRMs’ difficulty and redundancy cognition with a constructed heterogeneous dataset with inserted self-hypnosis.
-Experiments demonstrate over 70% length reduction on easy tasks and 40% on hard tasks with maintained performance stability, exhibiting clear difficulty-aware capabilities and reduced redundancy.
-Literature Review
-The advent of LRM (OpenAI et al.
-2024a; Guo et al.
-2025; Team 2025) enables breakthroughs in problem solving (Sun et al.
-2024b; Yu et al.
-2024) with Chain-of-Thought (Wei et al.
-2022; Wang et al.
-2023) and test-time scaling (Snell et al.
-2025).
-Nevertheless, this introduces the challenge of overthinking (Chen et al.
-2025; Team et al.
-2025).
-Works of efficient reasoning (Sui et al.
-2025) as summarized:
+To align with human System 1 and System 2 thinking patterns, we propose a novel two-stage fine-tuning strategy with a self-hypnosis mechanism for prospective and retrospective intervention, thus progressively inspiring LRMs’ difficulty and redundancy cognition.
+Experiments demonstrate over 70% length reduction on easy tasks and 40% on hard tasks with maintained robust performance, as shown in Fig.1, exhibiting clear difficulty-aware capabilities and reduced redundancy.
+2 Literature Review
+The advent of LRM OpenAI et al.
+(2024a); Guo et al.
+(2025); Team (2025) enables breakthroughs in problem solving Sun et al.
+(2024b); Yu et al.
+(2024) with Chain-of-Thought Wei et al.
+(2022); Wang et al.
+(2023) and test-time scaling Snell et al.
+(2025).
+Nevertheless, this introduces the challenge of overthinking Chen et al.
+(2025); Team et al.
+(2025); Sui et al.
+(2025).
+Prompt-based methods.
+Some works Aytes et al.
+(2025); Ding et al.
+(2024); Ning et al.
+(2024); Han et al.
+(2025); Xu et al.
+(2025); Lee et al.
+(2025) instruct LRMs by various prompts to respond concisely.
+Specifically, Renze and Guven (2024) explicitly prompts LLMs to perform concise step-by-step reasoning.
+SoT Ning et al.
+(2024) combines reasoning paradigms with linguistic constraints to minimize token usage.
+Token-Budget Han et al.
+(2025) establishes a self-estimated token budget as a constraint.
+Ge et al.
+(2025) regulates thinking token distribution by few-shot in-context learning.
+Output-based methods.
+Implicit-based methods Deng et al.
+(2024); Hao et al.
+(2024); Saunshi et al.
+(2025); Cheng and Durme (2024); Shen et al.
+(2025a) propose to internalize reasoning steps within the latent representation, thereby shifting the cognitive process from the explicit language space to the implicit hidden space.
+Saunshi et al.
+(2025) simulates multiple steps of CoT via a looped transformer layer to deepen the reasoning process internally.
+Another line of work focuses on dynamic execution.
+Sun et al.
+(2024a); Xie et al.
+(2023); Liao et al.
+(2025) optimize Best-of-N (BoN) decoding with speculation and rejection.
+Li et al.
+(2024); Yang et al.
+(2025a); Wang et al.
+(2025) explores potential early exit point.
+NoThinking Ma et al.
+(2025a) prompts reasoning models to directly output final answers without thinking.
 Model-based methods.
-Reinforcement Learning motivates some existing efficient reasoning works (Arora and Zanette 2025; Team et al.
-2025; Aggarwal and Welleck 2025) to incorporate length penalties.
-O1-pruner (Luo et al.
-2025b) introduces the Length-Harmonizing Reward, integrated with a PPO-style loss.
-DAST (Shen et al.
-2025b) constructs a length-preference dataset and employs SimPO (Meng and Xia 2024).
-AdaR1 (Luo et al.
-2025a) constructs a hybrid reasoning model by merging long and short CoT models and is optimized with DPO (Rafailov et al.
-2023).
-Apart from RL, which is vulnerable and expensive, fine-tuning is another effective post-training strategy (Yang et al.
-2025b).
-CoT-Value (Ma et al.
-2025b) regulates reasoning length by a controllable direction in parameter space, enabling dynamic length control.
-C3oT (Kang et al.
-2025) employs GPT-4 as a compressor while retaining key information.
-TokenSkip (Xia et al.
-2025) prunes less important tokens and fine-tune the model to skip the redundant tokens.
-Output-based methods.
-Implicit-based methods (Deng, Choi, and Shieber 2024; Hao et al.
-2024; Saunshi et al.
-2025; Cheng and Durme 2024; Shen et al.
-2025a) propose to internalize reasoning steps within the latent representation, transitioning reasoning from language space to hidden space.
-(Saunshi et al.
-2025) simulates multiple steps of CoT through a looped transformer layer, increasing the depth of reasoning without additional parameters.
-However, these implicit techniques introduce performance decline and affect interpretability.
-There are also dynamic criteria.
-For example, (Sun et al.
-2024a; Xie et al.
-2023; Liao et al.
-2025) optimize Best-of-N (BoN) decoding with speculation and rejection.
-(Li et al.
-2024; Yang et al.
-2025a; Wang et al.
-2025) explores possible early exit point.
-NoThinking (Ma et al.
-2025a) prompts reasoning models to directly output final answers without thinking, but heavily relies on sampling, which adversely increases overall overhead.
-Prompt-based methods.
-Some works (Aytes, Baek, and Hwang 2025; Ding et al.
-2024; Ning et al.
-2024; Han et al.
-2025; Xu et al.
-2025; Lee, Che, and Peng 2025) instruct LRMs by various prompts to respond concisely.
-Specifically, (Renze and Guven 2024) explicitly prompts LLMs to perform concise step-by-step reasoning.
-SoT (Ning et al.
-2024) combines reasoning paradigms with linguistic constraints to minimize token usage.
-Token-Budget (Han et al.
-2025) establishes a token budget estimated by LLM itself, and incorporates this estimation as a token number constraint.
-(Ge et al.
-2025) regulates thinking token distribution by few-shot in-context learning.
-Nevertheless, prompting relies on precise instruction following and robust task cognition, which will be proven unreliable later.
-Overall, most existing works treat efficient reasoning as a unified length compression paradigm.
-There is a necessity that has been widely overlooked that LRM may need the capacity to autonomously recognize and adapt to varying levels of problem difficulty, which aligns with System 1 and System 2 thinking patterns of human cognitive theory.
-Preliminary and Observation
-Background and Notations
+Fine-tuning offers a direct post-training strategy Yang et al.
+(2025b).
+CoT-Value Ma et al.
+(2025b) learns a controllable direction in the parameter space to steer generation length.
+C3oT Kang et al.
+(2025) employs GPT-4 as a compressor while retaining key information.
+TokenSkip Xia et al.
+(2025) fine-tunes the model to prune non-essential tokens during inference.
+Alternatively, reinforcement learning provides another vulnerable and expensive paradigm, often by incorporating length penalties to balance performance and efficiency Arora and Zanette (2025); Team et al.
+(2025); Aggarwal and Welleck (2025).
+O1-pruner Luo et al.
+(2025b) integrates a Length-Harmonizing Reward into a PPO-style loss.
+DAST Shen et al.
+(2025b) employs SimPO Meng and Xia (2024) on a constructed length-preference dataset, while AdaR1 Luo et al.
+(2025a) optimizes a merged hybrid long/short CoT model using DPO Rafailov et al.
+(2023).
+Limitations and our approach.
+Overall, existing approaches to efficient reasoning largely operate under a monolithic paradigm of length compression, which introduces critical limitations: 1) Prompt-based methods exhibit unreliability of instruction following and task cognition; 2) Implicit output-based methods suffer from performance degradation and compromised interpretability; 3) Model-based approaches are constrained by a behavioral rigidity that hinders adaptability.
+In contrast, we reframe overthinking not just as a behavioral flaw, but as a fundamental cognitive deficit, since length is not an accurate proxy for difficulty.
+Our approach instills a dual-level awareness: a global, prospective cognition of task difficulty to guide high-level strategy, and a local, retrospective in-progress perception of redundancy for fine-grained execution.
+This self-aware intervention mechanism cultivates the metacognitive skill of thinking about how to think efficiently, rather than simply enforcing brevity.
+3 Preliminary and Observation
+3.1 Background
 LLMs.
-Given a prompted question , conventional LLM generates a response by iteratively predicting each response token conditioned on and generated tokens .
-This is termed as System 1 (Li et al.
-2025b) thinking:
-where [Prompt] and [Answer] are question and response , which is often short and concise.
+Given a prompted question , LLM generates a response by iteratively predicting each response token conditioned on and generated tokens .
+This is termed as System 1 thinking Li et al.
+(2025b): , where [Prompt] and [Answer] are question and response , which is often short and concise.
 LRMs.
-Whereas in reasoning-focused LRMs, the generation process is split into two stages, e.g., thinking and conclusion, which is referred to as System 2 (Li et al. 2025b):
-where and are two delimiters denoting the start and end of reasoning, [Thoughts] and [Conclusion] are detailed reasoning process and concise answer, respectively.
-Formally, the reasoning part [Thoughts] reveals a trajectory of problem solving, known as CoT (Wei et al.
-2022; Wang et al.
-2023).
-At the micro level, some temporally consecutive tokens in [Thoughts] can be aggregated into short chunk based on semantics, logic and grammar:
-Similarly, each reasoning chunk in this paradigm is autoregressively generated by conditioning on question and previous reasoning thoughts: .
+Whereas in LRMs, the generation process is split into two stages, i.e., thinking and conclusion, referred to as System 2: , where and are two delimiters denoting the start and end of reasoning, [Thoughts] and [Conclusion] are detailed reasoning processes and concise answers, respectively.
+Formally, [Thoughts] reveals a trajectory of problem solving, known as CoT Wei et al.
+(2022); Wang et al.
+(2023).
+At the micro level, some consecutive tokens in [Thoughts] can be aggregated into short chunk based on semantics, logic and grammar:
+Similarly, each reasoning chunk is autoregressively generated by conditioning on question and previous chunks: .
 This explicit decomposition not only improves interpretability through transparent exposure of reasoning steps but also enhances the model’s capability for complex tasks.
-Specifically, some chunks are well structured by prefix keywords such as “Wait”, “However”, “Alternatively” (Guo et al.
-2025), marking the subsequent chunks will undergo logical backtracking, self-verification, or exploration.
-Works (Li et al.
-2025a) have shown that reflection structures are crucial to LRMs.
-Under this setting, the “overthinking” phenomenon is that many seem delicate yet contribute nothing to reasoning logic, resulting in prominent overhead.
-Observation: Reasoning Process Intervention
+Specifically, some reflective chunks are well structured by prefix keywords such as “Wait”, “However”, “Alternatively” Guo et al.
+(2025), marking the subsequent chunks will undergo logical backtracking, self-verification, or exploration, which have been proved crucial to LRMs Li et al.
+(2025a).
+Under this setting, we define “overthinking” as a phenomenon characterized by inefficiencies at both the global and local levels: globally, through an excessive number of reasoning steps () relative to task complexity, and locally, where individual steps () are elaborate yet logically superfluous, resulting in prominent overhead.
+3.2 Reasoning Process Intervention
 To address the overthinking issue, a straightforward approach is difficulty reminders—encouraging LRMs to self-regulate output length based on task difficulty judgment through prompts like “Please think quickly if you find it easy”.
-However, this conventional interventional technique necessitates two prerequisites: accurate instruction following and robust task difficulty cognition.
-While prior studies (Wu et al.
-2025; Han et al.
-2025) have highlighted the inherent challenges of the former, we have already shown that current LRMs struggle to recognize varying levels of problem difficulty in Fig.
+However, this hypothetical condition necessitates two prerequisites: accurate instruction following and robust task difficulty cognition.
+While prior studies Wu et al.
+(2025); Han et al.
+(2025) have highlighted the inherent challenges of the former, we have already shown that current LRMs lack cognition of problem difficulty in Fig.
 2.
-Instead of using a difficulty reminder in the prompt, we claim that a promising solution is transferring the intervention from the prompt to the inner reasoning process.
+Instead, we reveal that a promising solution is transferring the intervention from the prompt to the inner reasoning process.
 We conduct a pre-experiment in Fig.
-3, pretending the model has already correctly recognized the difficulty level of samples.
-Specifically, for a simple problem, we change the model input from “User: [Prompt]+[Question].
+3, pretending the model knows the difficulty level of samples.
+Given a simple problem, we change the model input from “User: [Prompt]+[Question].
 Assistant:” to “User: [Prompt]+[Question].
 Assistant: This is a simple question, let’s think quickly.”.
-While for hard problems, the corresponding predefined response prefix is “It seems difficult, let’s think thoroughly.”.
+While for hard problems, the predefined response prefix is “It seems difficult, let’s think thoroughly.”.
 Fig.
-3 shows that predefined difficulty awareness indeed helps to reduce response length, especially matched awareness (“simple, quickly” for easy and “difficult, thoroughly” for hard is matched, the reverse is unmatched).
-We define them as self-hypnosis that acts as a trigger in the output, guiding the subsequent inference trajectory.
-In fact, the reflection structures in LRMs (Guo et al.
-2025; Li et al.
-2025a) are essentially similar triggers aiming to intervene in subsequent think chunks.
-Compared to explicit prompts, we argue that these built-in triggers offer three key advantages: (1) native flexibility, (2) robustness, and (3) superior instruction-following capabilities.
-In the next section, we introduce how we guide LRMs to generate native interventional triggers of difficulty and redundancy cognition, autonomously.
-Methodology
-In this section, we describe our two-stage fine-tuning strategy Think-How-to-Think (TH2T) for mitigating overthinking, as shown in Fig. 4.
-Stage 1: Difficulty Cognition Injection
-We aim to endow LRM with the capability to assess the difficulty of a given problem.
+3 shows that predefined difficulty awareness largely outperforms reminder on reducing token length, especially matched awareness (“simple, quickly” for easy and “difficult, thoroughly” for hard is matched, the reverse is unmatched).
+We define them as self-hypnosis that acts as triggers in the output, guiding the subsequent inference trajectory.
+Compared to explicit prompts, these built-in triggers offer three key advantages: 1) native flexibility; 2) robustness; 3) superior instruction-following capabilities.
+In Sec.4, we utilize the self-hypnosis mechanism to guide LRMs to generate native interventional triggers of difficulty and redundancy cognition, autonomously.
+4 Methodology
+As discussed in Sec.3.1, the challenge of overthinking stems from failures in both high-level strategy selection and fine-grained step execution.
+We ground our method in the dual-process theory of cognition, i.e., intuitive System 1 and analytical System 2.
+The strategy calls for a global, prospective intervention applied pre-generation, while the execution demands a local, retrospective mechanism for in-progress correction and refinement.
+Since they are complementary and distinct systems, we propose a two-stage fine-tuning framework Think-How-to-Think (TH2T) for mitigating overthinking, as shown in Fig.
+4.
+4.1 Stage 1: Difficulty Cognition Injection
+The foundational principle of our approach is to cultivate the model’s capacity for task appraisal, which is a metacognitive prerequisite for adaptive reasoning.
 We define two datasets: for easy problems, and for difficult ones.
-Two models from the same series were selected to ensure consistency in response: for basic LLM with prompted Short-CoT capability (e.g., Qwen2.5 series) and for advanced LRM (e.g., DeepSeek-R1-Distill-Qwen2.5 series).
+Two models from the same series were selected to ensure consistency in response: for basic LLM with prompted Short-CoT capability and for advanced LRM.
 For each problem in , the responses are recorded as:
@@
 This motivates us to preserve ’s potential for efficiency on with satisfactory performance.
-Our desired thinking pattern of is responsed concisely on and elaborately on .
+The target reasoning pattern of entails concise responses on and detailed responses on .
 To align with this objective, we construct a draft dataset grounded in difficulty differentiation:
-where , similar to Eq. (5). aims to maintain the performance gain of on .
-To enable to perceive problem difficulty level autonomously, we introduce the trials of Fig.
-3.
-We inject specific difficulty-aware cognitive Hypnosis into the prefix of answer within , which, as demonstrated before, functions as interventional triggers guiding the reasoning trajectory:
+where . aims to maintain the performance gain of on .
+To enable to perceive problem difficulty level autonomously, we introduce the trials of Sec.
+3.2.
+We inject specific difficulty-aware cognitive Hypnosis into the prefix of answer within , which, as demonstrated before, act as an interventional signals for global, prospective strategy selection.
 If , the Difficulty Hypnosis H is defined as follows: “hypnosis This is a simple question, let’s think quickly./hypnosis”.
 If , then “hypnosisIt seems difficult, let’s think thoroughly./hypnosis”.
-The markers “hypnosis” and “/hypnosis” are defined as indicators for inserted difficulty hypnosis.
-The injected hypnosis is recognized as an embedded difficulty label to train to generate native difficulty cognition on the refined dataset , stimulating a concise thinking pattern on easy tasks.
-The execution of stage 1 is outlined at the upper section of the algorithmic framework Fig.
-(4).
-Stage 2: Redundancy Cognition Injection
-We extend difficulty-hypnosis of Stage 1 to redundancy-hypnosis, guiding to identify redundant structures within the reasoning process on hard tasks.
+The markers “hypnosis” and “/hypnosis” are indicators for inserted difficulty hypnosis.
+The injected hypnosis is recognized as an embedded difficulty label to train to generate native difficulty cognition, stimulating a concise thinking pattern on easy tasks.
+The execution of stage 1 is outlined in the upper section of Fig.
+4.
+4.2 Stage 2: Redundancy Cognition Injection
+We then extend difficulty-hypnosis to redundancy-hypnosis, guiding to identify redundant structures within the reasoning process on hard tasks.
 […]
-where is the difficulty-hypnosis introduced in Stage 1, are valid reasoning chunks to be kept, is the unnecessary reflection structure inside the reasoning process without effective contribution to the reasoning process.
-These chunks often carry out backtracking, self-verification, or exploration, but unfortunately, are nothing more than superfluous overhead.
-is a similar reflection structure; nonetheless, it primarily manifests in the final phase of the reasoning process, continuously iterating until attains the maximum output length.
+where is the difficulty-hypnosis introduced in Stage 1, are valid reasoning chunks to be kept.
+is the superfluous reflection structure inside the reasoning process without effective contribution to the reasoning process.
+is a similar reflection structure; nonetheless, it primarily manifests in the final phase of the reasoning, and the model enters an unproductive loop (trap), relentlessly iterating until reaches the maximum output length.
+This practice severely degrades inference efficiency, as latency often scales super-linearly with sequence length, and the vast majority of problems do not require such exhaustive thought.
 We define a Determinator to judge whether some is logically equivalent to or (i.e., redundant or looping).
@@
 In case of , original is refined as:
-wherein the redundancy hypnosis is defined as follows: “hypnosis Everything seems ok, let’s move on./hypnosis”. We use GPT-4 (OpenAI et al. 2024b) to implement the determinator.
-In case of , on one hand, the training set lacks a sufficient number of loop structures; on the other hand, excessively deferred looping hinders the generation of shorter-length samples.
+wherein the Redundancy Hypnosis is defined as follows: “hypnosis Everything seems ok, let’s move on./hypnosis”.
+We use GPT-4 OpenAI et al.
+(2024b) to implement the determinator.
+In case of , the training set lacks a sufficient number of loop structures; meanwhile, excessively deferred looping hinders the generation of shorter-length samples.
 Thus, we select intermediate chunk with words like “Wait”, “However”, “Alternatively” and iterate it as a simulation of looping, followed by hypnosis injection and truncation:
@@
 /hypnosis”.
-The choice of iterate number can be random, and we also try a probe-like approach (Zhang et al.
-2025).
-We use DeepSeek-R1 to identify chunks before which the prior content can already induce the correct answer, referred to as completeness requirements.
-The last-layer hidden states of the selected chunk tail is extracted to train a simple two-layer MLP as Determinator.
+To determine the optimal timing for truncation, we leverage DeepSeek-R1 to justify the earliest point—a condition we term the completeness requirement—where the preceding content becomes sufficient to derive the correct answer.
+The choice of iterate number can be random, and we also try a probe-like approach Zhang et al.
+(2025).
+The last-layer hidden states of the selected chunk tail are extracted to train a simple two-layer MLP as Determinator.
 Experiments indicate it yields similar results to the random approach, so we prefer the latter.
@@
 Meanwhile, generation continuity is ensured by remaining untruncated samples in the dataset, since we only apply Stage 2 to part of the samples in .
-Stage 2 execution is outlined in Fig.
-4.
-Experiments
-Experiment Setup
+5 Experiments
+5.1 Experiment Setup
 Models.
-Dataset Construction.
+We do experiments on Long-CoT models, including DeepSeek-R1-Distilled-Qwen-2.5-7B/14B/32B and -Llama-8B Guo et al.
+(2025).
+During data construction, the adopted Short-CoT model are Qwen2.5-7B/14B/32B Qwen et al.
+(2025) and Llama-3.1-8B-Instruct Grattafiori et al.
+(2024).
+They are and , respectively.
+Datasets.
 For dataset construction, we mainly adopt mathematical reasoning tasks, which are friendly for difficulty quantization.
@@
 In stage 2, redundancy and looping hypnosis are injected into difficult samples.
-The composition ratio of this heterogeneous dataset is 2:1:1.
-Baselines.
-We compare TH2T with two prompt-based methods, i.e., D-Prompt and NoThinking (Ma et al.
-2025a), and two fine-tuning methods, i.e., Tokenskip (Xia et al.
-2025), CoT-Valve (Ma et al.
-2025b).
-D-Prompt directly promotes LRM to regulate its output length based on self-judgment of task difficulty, through difficulty reminder in the prompt, as discussed in Preliminary Part.
-The remaining methods are from recent publications.
-| Methods | GSM8K | MATH-500 | ||||||||||
-| Acc | Gain | Len | Reduc | Latency | SpdUp | Acc | Gain | Len | Reduc | Latency | SpdUp | |
-| DeepSeek-R1-Distill-Qwen-7B | ||||||||||||
-| Original | 91.51 | 1057 | 20.29 | 85.2 | 2857 | 87.69 | ||||||
-| [1pt/2pt] D-Prompt | 90.59 | 0.92 | 906 | 14.3% | 16.38 | 1.24 | 86.0 | 0.8 | 2726 | 4.6% | 81.54 | 1.08 |
-| NoThinking | 88.47 | 3.04 | 243 | 76.9% | 3.34 | 6.07 | 79.4 | 5.8 | 699 | 75.5% | 11.81 | 6.43 |
-| [1pt/2pt] TokenSkip | 88.01 | 3.50 | 638 | 39.6% | 10.64 | 1.91 | 80.2 | 5.0 | 2109 | 26.2% | 53.94 | 1.62 |
-| CoT-Valve | 87.90 | 3.61 | 235 | 77.8% | 3.31 | 6.13 | 78.6 | 6.6 | 747 | 73.9% | 12.31 | 7.12 |
-| (w/o D.H.) | 90.83 | 0.68 | 503 | 47.4% | 7.80 | 2.60 | 85.2 | 0.0 | 1906 | 33.2% | 46.68 | 1.88 |
-| (w/o R.H.) | 91.96 | 0.48 | 397 | 62.4% | 5.99 | 3.39 | 86.0 | 0.8 | 2265 | 20.7% | 60.22 | 1.46 |
-| TH2T | 92.87 | 1.36 | 275 | 74.0% | 3.88 | 5.23 | 86.8 | 1.6 | 1772 | 38.0% | 41.65 | 2.11 |
-| DeepSeek-R1-Distill-Qwen-14B | ||||||||||||
-| Original | 93.63 | 584 | 9.38 | 86.6 | 2306 | 62.65 | ||||||
-| [1pt/2pt] D-Prompt | 93.91 | 0.28 | 606 | 3.9% | 9.95 | 0.94 | 86.6 | 0.0 | 2297 | 0.3% | 62.60 | 1.01 |
-| NoThinking | 88.41 | 3.48 | 206 | 64.9% | 3.06 | 3.07 | 76.0 | 10.6 | 612 | 73.5% | 10.05 | 6.23 |
-| [1pt/2pt] TokenSkip | 89.30 | 4.33 | 314 | 46.2% | 4.79 | 1.96 | 73.2 | 13.4 | 1356 | 41.2% | 28.11 | 2.23 |
-| CoT-Valve | 92.70 | 0.90 | 526 | 9.9% | 8.13 | 1.15 | 84.2 | 2.4 | 2111 | 8.4% | 53.95 | 1.16 |
-| (w/o D.H.) | 93.69 | 0.06 | 338 | 42.2% | 4.82 | 1.95 | 86.2 | 0.4 | 1549 | 32.8% | 33.93 | 1.85 |
-| (w/o R.H.) | 93.47 | 0.16 | 299 | 51.2% | 4.34 | 2.16 | 85.8 | 0.8 | 1706 | 26.0% | 39.83 | 1.57 |
-| TH2T | 94.08 | 0.45 | 246 | 58.1% | 3.38 | 2.78 | 86.4 | 0.2 | 1343 | 41.8% | 27.84 | 2.25 |
-| DeepSeek-R1-Distill-Qwen-32B | ||||||||||||
-| Original | 95.60 | 717 | 12.14 | 87.2 | 2357 | 63.02 | ||||||
-| [1pt/2pt] D-Prompt | 95.60 | 0.00 | 663 | 7.6% | 11.12 | 1.09 | 86.8 | 0.4 | 2198 | 6.8% | 57.11 | 1.10 |
-| NoThinking | 93.86 | 1.74 | 228 | 68.2% | 3.25 | 3.74 | 80.2 | 7.0 | 659 | 72.0% | 10.98 | 5.74 |
-| [1pt/2pt] TokenSkip | 96.54 | 0.94 | 423 | 41.0% | 6.28 | 1.93 | 79.8 | 7.4 | 1567 | 33.5% | 34.26 | 1.84 |
-| CoT-Valve | 96.10 | 0.50 | 781 | 8.9% | 12.56 | 0.97 | 85.3 | 1.9 | 2263 | 3.9% | 60.21 | 1.05 |
-| (w/o D.H.) | 96.08 | 0.48 | 361 | 49.7% | 5.47 | 2.22 | 87.4 | 0.2 | 1863 | 20.9% | 45.18 | 1.39 |
-| (w/o R.H.) | 95.75 | 0.15 | 323 | 55.0% | 4.84 | 2.51 | 86.8 | 0.4 | 2114 | 10.3% | 53.99 | 1.17 |
-| TH2T | 97.04 | 1.44 | 263 | 63.3% | 3.51 | 3.46 | 88.0 | 0.8 | 1733 | 26.5% | 40.13 | 1.57 |
-| Bench | Difficulty Level | Cognition | |
-| Original | TH2T | ||
-| GSM8K | all easy | 14.9% | 98.3% |
-| MATH | lv.1-2 (easy) | 9.8% | 93.4% |
-| lv.4-5 (hard) | 12.0% | 91.5% | |
-| AIME2024 | all hard | 16.7% | 96.7% |
-| OmniMath | lv.1-2 (easy) | 20.0% | 96.7% |
-| lv.4-5 (hard) | 13.3% | 93.3% | |
-Evaluations.
-We evaluate on 4 mathematical benchmarks (GSM8K, MATH, AIME2024 and Omni-MATH) and 1 QA benchmark GPQA (Rein et al.
-2024).
+The composition ratio of this heterogeneous dataset is 2:1:1, yielding samples.
+We evaluate on 4 mathematical tasks (GSM8K, MATH, AIME2024 and Omni-MATH) and 1 QA benchmark GPQA Rein et al.
+(2024).
 Metrics include accuracy, average number of output tokens, and inference latency per sample.
 We also report the success rate of the difficulty awareness, redundancy, and looping optimization.
-The detailed experiment settings are in the appendix.
-Experimental Results
-Main Results.
-The performance of TH2T on GSM8K and MATH is illustrated in Tab.
+Baselines.
+We compare TH2T with the following methods: (1) D-Prompt (prompt-based) directly regulates response length based on self-judgment of task difficulty via the Difficulty Reminder in Prompt (in Sec.3.2).
+(2) NoThinking Ma et al.
+(2025a) (output-based) bypasses the thinking process via prompting for all difficulty-level questions.
+(3) Tokenskip Xia et al.
+(2025) (model-based) fine-tunes the model to prune non-essential tokens during inference, regardless of task difficulty.
+(4) CoT-Valve Ma et al.
+(2025b) (model-based) learns a controllable direction in the parameter space to steer generation length based on task difficulty.
+(5) AdaCtrl111We re-implemented all baselines except AdaCtrl (citing its original 7B results due to reproduction complexity).
+Huang et al.
+(2025) (model-based) uses reinforcement learning and allocates reasoning budgets based on self-assessed problem difficulty, which is more of a label-driven distillation.
+| Methods | GSM8K | MATH-500 | AIME2024 | OmniMath | GPQA | ||||||||||
+| Acc. | Len. | Reduc. | Acc. | Len. | Reduc. | Acc. | Len. | Reduc. | Acc. | Len. | Reduc. | Acc. | Len. | Reduc. | |
+| DeepSeek-R1-Distill-Qwen-7B | |||||||||||||||
+| Original | 91.51 | 1057 | 85.2 | 2857 | 50.0 | 10570 | 45.0 | 5736 | 32.8 | 5349 | |||||
+| [1pt/2pt] D-Prompt | 90.59 | 906 | 14.3% | 86.0 | 2726 | 4.6% | 46.7 | 10209 | 3.4% | 45.0 | 5398 | 5.9% | 30.3 | 5117 | 4.3% |
+| NoThinking | 88.47 | 243 | 76.9% | 79.4 | 699 | 75.5% | 40.0 | 4134 | 60.9% | 40.0 | 2083 | 63.7% | 29.3 | 2084 | 61.0% |
+| TokenSkip | 88.01 | 638 | 39.6% | 80.2 | 2109 | 26.2% | 40.0 | 5559 | 47.4% | 38.3 | 3461 | 39.7% | 28.8 | 4001 | 25.2% |
+| CoT-Valve | 87.90 | 235 | 77.8% | 78.6 | 747 | 73.9% | 43.3 | 5871 | 44.5% | 41.7 | 3311 | 42.3% | 28.3 | 3226 | 39.7% |
+| AdaCtrl | 90.98 | 349 | - | 74.0 | 3196 | - | 21.3 | 16889 | - | - | - | - | - | - | - |
+| [1pt/2pt] (w/o D.H.) | 90.83 | 503 | 47.4% | 85.2 | 1906 | 33.2% | 50.0 | 8213 | 22.3% | 46.7 | 3571 | 37.7% | 33.3 | 4372 | 18.3% |
+| (w/o R.H.) | 91.96 | 397 | 62.4% | 86.0 | 2265 | 20.7% | 46.7 | 8684 | 17.8% | 48.3 | 4247 | 26.0% | 32.8 | 4481 | 16.2% |
+| TH2T | 92.87 | 275 | 74.0% | 86.8 | 1772 | 38.0% | 50.0 | 7490 | 29.1% | 48.3 | 3377 | 41.1% | 32.8 | 4166 | 22.1% |
+| DeepSeek-R1-Distill-Qwen-14B | |||||||||||||||
+| Original | 93.63 | 584 | 86.6 | 2306 | 50.0 | 9900 | 46.7 | 5373 | 50.0 | 3847 | |||||
+| [1pt/2pt] D-Prompt | 93.91 | 606 | 3.9% | 86.6 | 2297 | 0.3% | 50.0 | 10408 | 5.1% | 41.7 | 4994 | 7.1% | 49.0 | 3946 | 2.6% |
+| NoThinking | 88.41 | 206 | 64.9% | 76.0 | 612 | 73.5% | 43.3 | 3735 | 62.3% | 43.3 | 1944 | 63.8% | 45.5 | 1454 | 62.2% |
+| TokenSkip | 89.30 | 314 | 46.2% | 73.2 | 1356 | 41.2% | 46.7 | 6772 | 31.6% | 40.0 | 3158 | 41.2% | 46.9 | 2465 | 35.9% |
+| CoT-Valve | 92.70 | 526 | 9.9% | 84.2 | 2111 | 8.4% | 40.0 | 7027 | 29.0% | 41.7 | 2751 | 48.8% | 46.5 | 2872 | 25.3% |
+| [1pt/2pt] (w/o D.H.) | 93.69 | 338 | 42.2% | 86.2 | 1549 | 32.8% | 46.7 | 7807 | 21.1% | 46.7 | 3745 | 30.3% | 50.5 | 2208 | 42.6% |
+| (w/o R.H.) | 93.47 | 299 | 51.2% | 85.8 | 1706 | 26.0% | 50.0 | 8156 | 17.6% | 48.3 | 3858 | 28.2% | 50.5 | 2463 | 36.0% |
+| TH2T | 94.08 | 246 | 58.1% | 86.4 | 1343 | 41.8% | 50.0 | 7140 | 27.9% | 48.3 | 3486 | 35.1% | 51.0 | 1901 | 50.6% |
+| DeepSeek-R1-Distill-Qwen-32B | |||||||||||||||
+| Original | 95.60 | 717 | 87.2 | 2357 | 60.0 | 9605 | 50.0 | 5540 | 52.0 | 4384 | |||||
+| [1pt/2pt] D-Prompt | 95.60 | 663 | 7.6% | 86.8 | 2198 | 6.8% | 56.7 | 9445 | 1.7% | 48.3 | 5135 | 7.3% | 49.5 | 3877 | 11.6% |
+| NoThinking | 93.86 | 228 | 68.2% | 80.2 | 659 | 72.0% | 53.3 | 3208 | 66.6% | 45.0 | 2146 | 61.3% | 50.5 | 1627 | 62.9% |
+| TokenSkip | 96.54 | 423 | 41.0% | 79.8 | 1567 | 33.5% | 50.0 | 5856 | 39.0% | 43.3 | 3730 | 32.7% | 47.4 | 2874 | 34.4% |
+| CoT-Valve | 96.10 | 781 | 8.9% | 85.3 | 2263 | 3.9% | 53.3 | 7997 | 16.7% | 43.3 | 4663 | 15.8% | 49.0 | 3550 | 19.0% |
+| [1pt/2pt] (w/o D.H.) | 96.08 | 361 | 49.7% | 87.4 | 1863 | 20.9% | 56.7 | 7737 | 19.4% | 48.3 | 3432 | 38.1% | 51.5 | 2513 | 42.7% |
+| (w/o R.H.) | 95.75 | 323 | 55.0% | 86.8 | 2114 | 10.3% | 60.0 | 8155 | 15.1% | 45.0 | 3644 | 34.2% | 51.0 | 2809 | 35.9% |
+| TH2T | 97.04 | 263 | 63.3% | 88.0 | 1733 | 26.5% | 56.7 | 7004 | 27.1% | 48.3 | 2998 | 45.9% | 52.0 | 2367 | 46.0% |
+| Bench | Difficulty lv. | Cognition Acc. | Confidence | ||
+| Original | TH2T | Original | TH2T | ||
+| GSM8K | all easy | 14.9% | 98.3% | 15.46 | 17.19 |
+| MATH | lv.1-2 (easy) | 9.8% | 93.4% | 16.09 | 16.97 |
+| lv.4-5 (hard) | 12.0% | 91.5% | 15.72 | 16.25 | |
+| AIME2024 | all hard | 16.7% | 96.7% | 16.64 | 16.89 |
+| OmniMath | lv.1-2 (easy) | 20.0% | 96.7% | 16.41 | 17.05 |
+| lv.4-5 (hard) | 13.3% | 93.3% | 16.94 | 17.76 | |
+5.2 Results
+TH2T achieves a significant reduction in computational overhead without compromising performance.
+The in-distribution performance on GSM8K and MATH is illustrated in Tab.
 1.
@@
 2) TH2T achieves the best reduction in reasoning length among methods with stable accuracy—up to 74% on GSM8K and 38% on MATH with the 7B model, translating to over 5 and 2 latency speedups, respectively.
-3) Greater gains on easier tasks, validating TH2T’s ability to adapt reasoning effort based on task difficulty.
+3) Greater gains are achieved on easier tasks, validating TH2T’s ability to adapt reasoning effort based on task difficulty.
 4) These observations are consistent across 7B, 14B and 32B models.
 As shown in Fig.
-1, other methods (e.g., NoThinking and CoT-Value) achieve comparable compression ratios in token length but at the cost of severe accuracy degradation (often 10 points), underscoring TH2T’s robustness in balancing efficiency and reliability.
-Difficulty Cognition.
-Does TH2T present better difficulty cognition?.
-As shown in Tab.
-2, the entirety of GSM8K, MATH lv.1-2, and Omni-Math lv.1-2 are set as “Easy” while the remaining are set as “Hard”.
-The original R1-Distilled-7B model is prompted to predict its own understanding of problem difficulty and choose between “Easy”, “Medium”, and “Hard”.
-TH2T can response with native self-hypnosis in the prefix like “This is a simple question” or “It seems difficult”, which indicates the cognition of task difficulty as “Easy” or “Hard”.
+1, other methods (e.g., NoThinking and CoT-Value) achieve comparable compression ratios in token length but at the cost of severe accuracy degradation (often 10 points), underscoring TH2T’s robustness in balanced efficiency.
+TH2T substantially strengthens the calibration of task difficulty cognition.
+The entirety of GSM8K, MATH lv.1-2, and Omni-Math lv.1-2 are set as “Easy” while the remaining are “Hard”.
 As shown in Tab.
 2 and Fig.
-2, the original LRM demonstrates ambiguous recognition of task difficulty, uniformly categorizing tasks of varying complexity as medium level.
-Whereas, our approach exhibits a clear distinction with high precision over 90+%, validating the effectiveness of the first stage of TH2T.
-Redundancy Cognition.
-Does TH2T present better redundancy cognition?
+2, the original LRM reveals ambiguous recognition of task difficulty (typically ), defaulting to the Medium label in the 3-option setting and lacking discriminability in the 2-option setting.
+In contrast, our approach yields a clear distinction with precision above 90%, evidencing effective calibration of task difficulty cognition enabled by the first-stage training.
+TH2T substantially enhances cognition of redundancy.
 Tab.
-3 shows the average number of reflective and looping structures.
-Comparison is made between R1-Distilled-7B and Ours-7B.
-We observe that: 1) TH2T effectively suppresses the number of reflection chunks by 10 on easy tasks and 3 on hard tasks, suggesting that extensive reflection is unnecessary for strong reasoning.
-2) The looping problem in TH2T undergoes substantial restraint.
-Notably, 59.5% of incorrect answers by the original model on MATH exhibit tail recursion, which is mitigated to 21.4% in ours.
+3 reports the average number of reflective and looping structures, comparing R1-Distilled-7B and TH2T-7B.
+We observe that: 1) TH2T effectively suppresses the number of reflection chunks by 10 on easy tasks and 3 on hard tasks, suggesting that extensive reflection is superfluous for strong reasoning.
+2) TH2T substantially alleviates the looping problem.
+Notably, 59.5% of incorrect answers by the original model on MATH contain tail recursion, which is reduced to 21.4% under TH2T.
 As shown in Fig.
-6(b), we observe zero loop structure in GSM8K.
-This distinct pattern of short-length concentration underscores the efficacy of Stage-2 of TH2T, presenting a more concise reasoning structure.
+5(b), we observe zero loop structure in GSM8K.
+This distinct pattern of short-length concentration highlights the efficacy of Stage-2 of TH2T, yielding more concise and efficient reasoning structures.
 | Bench | Split | Reflective | Looping | ||
@@
 | –incorrect | 96.5 | 39.1 | 59.5% | 21.4% | |
+5.3 Analysis
 Generalization Ability.
-To validate the generalization on OOD tasks, we evaluate on 2 more mathematical benchmarks and 1 QA benchmark. As shown in Fig. 5, accuracy maintenance and token length reduction are consistent.
-Results of Llama series model are in the appendix.
-Ablation Study
-Ablation on Stages.
+To validate the generalization on OOD tasks, we evaluate on 2 more mathematical benchmarks, i.e., AIME2024 and OmniMath, and 1 QA benchmark, i.e., GPQA.
+As shown in Tab.
+1, consistent with the main results, our approach preserves accuracy while yielding substantial token reduction up to 20-50%.
+Internal states validate the enhanced cognition.
+To examine whether our method instills a genuine sense of difficulty cognition, we investigate the model’s internal state in Tab.2.
+We posit that the confidence Fu et al.
+(2025) of the first token acts as an intrinsic, metacognitive signal of the model’s perceived task difficulty.
+Our method exhibits higher confidence, aligning with a concurrent rise in difficulty assessment accuracy.
+This consistency serves as a clear, self-consistent internal signal, providing strong validation that the enhanced cognition is explicitly represented in internal state.
+Ablation on two-stage design.
 We explore two more settings: 1) without Difficulty Hypnosis in Stage-1, and 2) without Redundancy Hypnosis in Stage-2, which are annotated as w/o D.H.
@@
 1.
-From the perspective of response reduction, we observe that D.H.
-introduces more gain on easy task GSM8K than on hard task MATH (e.g., 26.6% vs 4.8%, on 7B model), while R.H.
+We observe that D.H.
+introduces more token reduction on easy task GSM8K than on hard task MATH (e.g., 26.6% vs 4.8%, on 7B model), while R.H.
 exerts a greater impact on MATH (11.6% vs 17.3%).
-As we illustrate during method formulation, we design Difficulty Cognition Injection Stage to build native perception on easy tasks and Redundancy Cognition Injection Stage to mitigate non-essential structures that frequently occur in hard tasks.
-These results demonstrate the rationality of our two-stage strategy.
-Ablation on Starting Model.
-Instead of training on LRMs, we also try directly starting from a Short-CoT model with the same data in Tab.
-4, which reveals a notable failure in performance and response length.
-This motivates us to include an underlying Stage-0 in Fig.
-4 to start from an LRM, instead of teaching a Short-CoT model from scratch.
-| Method | GSM8K | MATH-500 | ||
-| Accuracy | Length | Accuracy | Length | |
-| Qwen2.5-7B | 86.20 | 280 | 62.80 | 629 |
-| [1pt/2pt] TH2T | 72.48 | 238 | 66.20 | 2333 |
-| R1-Distill-7B | 91.51 | 1057 | 85.2 | 2857 |
-| [1pt/2pt] TH2T | 92.87 | 275 | 86.8 | 1772 |
-Ablation on Training Steps.
-We examine the dynamics of accuracy and average token length during the training process (figure in appendix), and obtain two interesting findings: 1) As training steps increase, accuracy initially declines, then returns to the original level, with occasional positive gains during the process.
+As we illustrate in Sec.4, we design Difficulty Cognition Injection Stage to build native perception on easy tasks and the Redundancy Cognition Injection Stage to mitigate non-essential structures that frequently occur in hard tasks.
+This phenomenon also holds for the other three datasets, demonstrating the rationality of our two-stage framework.
+Acc and length across difficulty levels.
+We report a breakdown of the accuracy and average token length distribution across the complete spectrum of difficulty levels on 4 tasks in Fig.6.
+We compare the original R1-Distilled-Qwen-7B with TH2T-7B, where higher numerical levels indicate greater question difficulty.
+Our observations are as follows: 1) Both models produce progressively longer and more elaborate responses with increasing difficulty.
+2) Across all difficulty levels and tasks, TH2T consistently reduces response length while maintaining stable performance.
+3) Length reduction is most pronounced for easier questions.
+6 Conclusion
+This paper presents TH2T, a novel two-stage fine-tuning strategy for large reasoning models (LRMs) that effectively addresses the challenge of overthinking.
+Our empirical analysis indicates that current LRMs are largely limited to recognizing task difficulty, exhibiting a necessity for aligning with human System 1 and System 2 thinking patterns.
+Motivated by this, TH2T progressively enables native difficulty and redundancy cognition based on a hybrid dataset and proposed self-hypnosis mechanism.
+To the best of our knowledge, we are the first to inject difficulty/redundancy hypnosis as built-in triggers for global, prospective, or local, retrospective interventional signals.
+Experiments demonstrate that TH2T significantly reduces inference length without compromising performance, exhibiting clear sensitivity to difficulty and redundancy.
+Limitations
+In real-world cognitive settings, difficulty exists on a continuum, unlike the 2- or 3-level discretization employed in this study.
+However, to maintain research focus and interpretability, we adopt this simplified cognitive modeling approach.
+Also, our current approach primarily focuses on standard fine-tuning using the generated hybrid dataset.
+While this method is effective within our current framework, it may not fully exploit the potential for optimized efficiency.
+Reinforcement learning presents a promising avenue for further performance gains, although it can be unstable and computationally demanding.
+Finally, we acknowledge that, in theory, models of parameter scales exhibit distinct capability boundaries, which result in varying levels of task difficulty perception.
+At this stage, we apply a unified difficulty evaluation standard across models of different scales.
+Exploring the relationship between a model’s cognition and its capability boundaries is a promising research direction, potentially by dynamic correctness ratio during sampling, despite its significant computational overhead.
+Appendices
+| Methods | GSM8K | MATH-500 | AIME2024 | OmniMath | GPQA | Average | ||||||
+| Acc | Len | Acc | Len | Acc | Len | Acc | Len | Acc | Len | Acc | Len | |
+| DeepSeek-R1-Distill-Llama-8B | ||||||||||||
+| Original | 75.36 | 260 | 61.8 | 701 | 3.3 | 1378 | 26.67 | 844 | 34.34 | 704 | 40.29 | 777 |
+| D-Prompt | 75.14 | 261 | 62.2 | 596 | 6.6 | 1100 | 28.33 | 808 | 34.34 | 669 | 41.32 | 687 |
+| NoThinking | 75.24 | 257 | 62.2 | 580 | 6.6 | 1019 | 28.33 | 714 | 32.32 | 619 | 40.94 | 638 |
+| TH2T | 75.58 | 245 | 61.6 | 498 | 13.3 | 861 | 40.00 | 561 | 40.91 | 526 | 46.28 | 538 |
+Appendix A More Experimental Results
+A.1 Results on Llama Model
+We provide detailed experimental results of the R1-Distilled-Llama-8B model evaluated on 4 mathematical benchmarks (i.e., GSM8K, MATH, AIME2024, and OmniMath) and 1 QA benchmark (i.e., GPQA).
+We report accuracy and average token length for comparison in Tab.
+4.
+A.2 Ablation on Training Steps.
+We examine the dynamics of accuracy and average token length during the training process (in Fig.
+7), and obtain two interesting findings: 1) As training steps increase, accuracy initially declines, then returns to the original level, with occasional positive gains during the process.
 2) For hard tasks like MATH, the response length exhibits a reverse increase after exceeding a certain number of steps.
 This inspires us to set the training steps to 6400 to prevent model degradation.
-Conclusion
-This paper presents TH2T, a novel two-stage fine-tuning strategy for large reasoning models (LRMs) that effectively addresses the challenge of overthinking.
-Our empirical analysis reveals that current LRMs are primarily limited to recognizing task difficulty, exhibiting a necessity for aligning with human System 1 and System 2 thinking patterns.
-Motivated by this, TH2T progressively enables native difficulty and redundancy cognition based on a heterogeneous dataset and proposed self-hypnosis mechanism.
-To the best of our knowledge, we are the first to inject difficulty/redundancy/looping hypnosises as built-in triggers for reasoning trajectory intervention.
-Experiments demonstrate that TH2T significantly reduces inference length with maintained performance stability, exhibiting clear difficulty and redundancy sensitivity.
-Future research can explore finer-grained reasoning trajectory intervention.
-Supplementary Materials
-More Experimental Details
-Dataset Construction
+A.3 Visualization of the Effect on Token Length
+We collect the first 50 samples from GSM8K and MATH test set in sequential order, and compare the output token length distribution between our TH2T and the original DeepSeek-R1-Distill-Qwen-7B in Fig.
+8.
+We can see TH2T significantly reduces response length.
+Appendix B More Experimental Details
+B.1 Dataset Construction
 The datasets used in this paper incorporate difficulty level labeling as a key consideration, so we mainly use mathematical reasoning tasks.
@@
 Meanwhile, the redundancy hypnosis and looping hypnosis are also injected into in a ratio of 1:1.
-The overall composition is set as 2:1:1.
-Baselines
-In our main experiments, we compare TH2T with two prompt-based methods and two fine-tuning methods.
-D-Prompt directly promotes LRM to regulate its output length based on self-judgment of task difficulty based on Difficulty Reminder in Prompt, as discussed in our Preliminary Part.
-NoThinking bypasses the thinking process via prompting for all difficulty-level questions.
-Tokenskip constructs a dataset where less important tokens are pruned and fine-tuned to enable the model to skip the redundant tokens in inference, regardless of task difficulty.
-CoT-Valve enables dynamic control of CoT length by identifying and leveraging a controllable direction in parameter space, reducing inference cost by compressing reasoning paths based on task difficulty.
-We replicate the first two baselines and report the relative performance improvement of the latter two from the raw papers due to variability of the model and implementation methods.
-Benchmarks
+The overall composition is set as 2:1:1, yielding samples in total.
+B.2 Benchmarks
 We evaluate on 4 mathematical benchmarks and 1 QA benchmark.
@@
 To explore generalization ability in other domains, we include the GPQA benchmark to test performance on QA tasks.
-Metrics
-Main results are assessed on three metrics: accuracy, average number of output tokens, and inference latency per sample.
+B.3 Metrics
+Main results are assessed on accuracy and the average number of output tokens.
 To substantiate the research objectives, we incorporate the success rate of the difficulty awareness.
 We also report the redundancy cognition optimization, which includes the reduction of redundant structures (e.g., “Wait”, “Alternatively”, or “Check”), as well as the reflection looping ratio.
-Implementation Details
+To examine whether our method instills a genuine sense of difficulty cognition, we investigate the model’s internal state, i.e., the confidence value of the first generated token.
+B.4 Implementation Details
 We adopt LoRA, a widespread LLM fine-tuning approach, to train our models with a rank , scaling parameter , and learning rate .
-All experiments are conducted on 2NVIDIA H800 GPU (80GB).
-For the decoding strategy, we employ greedy decoding, and the maximum generation length equals 8192 (except 16,384 for AIME2024).
+All experiments are conducted on 4NVIDIA H800 GPU (80GB).
+For the decoding strategy, we employ greedy decoding.
+The maximum generation length is set to 8192 for GSM8K and MATH500, and 16,384 for AIME2024, OmniMath, and GPQA.
 The random seed is fixed to 42.
-| Methods | GSM8K | MATH-500 | AIME2024 | OmniMath | GPQA | Average | ||||||
-| Acc | Len | Acc | Len | Acc | Len | Acc | Len | Acc | Len | Acc | Len | |
-| DeepSeek-R1-Distill-Llama-8B | ||||||||||||
-| Original | 75.36 | 260 | 61.8 | 701 | 3.3 | 1378 | 26.67 | 844 | 34.34 | 704 | 40.29 | 777 |
-| D-Prompt | 75.14 | 261 | 62.2 | 596 | 6.6 | 1100 | 28.33 | 808 | 34.34 | 669 | 41.32 | 687 |
-| NoThinking | 75.24 | 257 | 62.2 | 580 | 6.6 | 1019 | 28.33 | 714 | 32.32 | 619 | 40.94 | 638 |
-| TH2T | 75.58 | 245 | 61.6 | 498 | 13.3 | 861 | 40.00 | 561 | 40.91 | 526 | 46.28 | 538 |
-More Experimental Results
-Results on Llama Model
-We provide detailed experimental results of the R1-Distilled-Llama-8B model evaluated on 4 mathematical benchmarks (i.e., GSM8K, MATH, AIME2024, and OmniMath) and 1 QA benchmark (i.e., GPQA).
-We report accuracy and average token length for comparison in Tab.
-5.
-Acc and Length on Detailed Difficulty Levels
-We report a breakdown of the accuracy and average token length distribution across the complete spectrum of difficulty levels on 4 tasks.
-The comparison is made between original R1-Ditstilled-Qwen–7B and TH2T-7B.
-Each question in the MATH and OmniMath datasets is labeled with a difficulty level, ranging from 1 to 5 and from 1 to 8, respectively.
-The higher the numerical level, the more difficult the question.
-Meanwhile, all questions in GSM8K and AIME2024 are considered easy and very difficult, respectively.
-Results are illustrated in Fig.7, and we obtain the following observations: 1) Both models generate progressively longer and more elaborate responses as the difficulty level increases.
-2) across all difficulty levels in all tasks, TH2T exhibits significant length reduction, as well as stable performance maintenance.
-3) The lower the difficulty level, the more significant the reduction in length.
-Dynamics of Accuracy and Token Length
-We report the dynamics of accuracy and average token length during the training steps in Fig.
-8, and obtain two interesting findings: 1) As training steps increase, accuracy initially declines, then returns to the original level, with occasional positive gains during the process.
-2) For hard tasks like MATH, the response length exhibits a reverse increase after exceeding a certain number of steps.
-This inspires us to set the training steps to 6400 to prevent model degradation.
-Visualization of the Effect on Token Length
-We collect the first 50 samples from GSM8K and MATH test set in sequential order, and compare the output token length distribution between our TH2T and the original DeepSeek-R1-Distill-Qwen-7B in Fig.
-9.
-We can see TH2T significantly reduces response length.
-More Explanation on Motivation
+Appendix C More Explanation on Motivation
 We investigate the efficiency issue from a suspicion: Are current LRMs treating various task types equally with identical response strategy?
@@
 From Fig.
-10, we obtain three findings:
+9, we obtain three findings:
 Across the easy and difficult tasks, the current reasoning model shows similar token length scaling ratio (i.e., both around 400%), which verifies our initial suspicion that LRMs are allocating similar additional computing resources to tasks with varying difficulty levels.
-Despite incurring minimal costs, basic LLM can still replicate the majority of correct responses (89.9%) of LRM’s on the easy tasks.
+Despite incurring minimal costs, basic LLM can still replicate the majority of correct responses (89.9%) of LRM’s on the easy tasks, which is why we claimed that can replicate the majority correct responses of on in Sec.4.1.
 Compared with difficult tasks, the easy task only witnesses a minor accuracy gain (5.3% vs 22.4%) under the same computation resource allocation strategy, which is not what we expected.
@@
 These raise a reasonable speculation: LRMs fall short of perceiving task complexity, since massive inference resources on easy ones are superfluous and lead to limited performance gain.
-As discussed in Results on Redundancy Cognition of the main paper, we observe severe redundant structures in LRM’s responses to hard tasks, including unnecessary reflection along with a remarkably high proportion of reasoning loops in incorrect answers.
+As discussed in Results on Redundancy Cognition of the main paper, we observe severe redundant structures in LRM’s responses to hard tasks, including superfluous reflection along with a remarkably high proportion of reasoning loops in incorrect answers.
 These findings lead to the two objectives of this paper:
@@
 Mitigating LRM overthinking by promoting redundancy cognition of internal reasoning structure (results and analysis are presented in the Results of Redundancy Cognition section of the main paper).
-Prompts
+Appendix D Prompts
 For reproducibility, we provide the complete list of prompts adopted in our experiments.
-Difficulty level evaluation
+D.1 Difficulty level evaluation
 Fig. 10 shows the prompt template we used during difficulty level evaluation.
-Baseline
+D.2 Baseline
 Fig. 11 is the difficulty reminder for the baseline D-Prompt, which is also adopted in the pre-experiment. Fig. 12 is the setting of baseline NoThinking.
-Design of Determinator
+Appendix E Design of Determinator
 During Stage 2 of TH2T, we design a Determinator.
@@
 14 and Fig.
-16.
+15.
 Furthermore, when injecting hypnosis, the completeness requirement in Fig.
 13 is adopted.
-Limitations and Future Work
-Our current approach primarily centers on standard fine-tuning using our generated heterogeneous dataset.
-While this method effectively enhances model performance within our present framework, it may not fully leverage the potential for optimized efficiency.
-Advanced reinforcement learning (RL)-based training strategies may offer a promising avenue for further performance improvements.
-However, RL can be very unstable and computationally intensive.
-Due to computational constraints, we defer the exploration of sophisticated RL techniques to future research.
-We also recognize that theoretically, models of different series and parameter scales have distinct capability boundaries, leading to varying levels of task difficulty perception.
-In this phase, we introduce a common difficulty evaluation standard applicable to models at different scales.
-We plan to explore difficulty definition methods tailored to models of variant capability ranges in the future, aiming to advance more autonomous difficulty adaptation and perception strategies.
-One possible solution is to depend on the correctness ratio during sampling.
-However, a significant limitation of this method lies in its substantially increased computational overhead.
-Case Study
+Appendix F Case Study